### v3

# Reliability-Latency-Rate Tradeoff in Low-Latency Communications with Finite-Blocklength Coding

Abstract
Low-latency communication plays an increasingly important role in delay-sensitive applications by ensuring the real-time information exchange. However, due to the constraint on the maximum instantaneous power, guaranteeing bounded latency is challenging. In this paper, we investigate the reliability-latency-rate tradeoff in low-latency communication systems with finite-blocklength coding (FBC). Specifically, we are interested in the fundamental tradeoff between error probability, delay-violation probability (DVP), and service rate. Based on the effective capacity (EC), we present the gain-conservation equations to characterize the reliability-latency-rate tradeoffs in low-latency communication systems. In particular, we investigate the low-latency transmissions over an additive white Gaussian noise (AWGN) channel and a Nakagami-$m$ fading channel. By defining the service rate gain, reliability gain, and real-time gain, we conduct an asymptotic analysis to reveal the fundamental reliability-latency-rate tradeoff of ultra-reliable and low-latency communications in the high signal-to-noise-ratio (SNR) regime. To analytically evaluate and optimize the quality-of-service-constrained throughput of low-latency communication systems adopting FBC, an EC-approximation method is conceived to derive the closed-form expression of that throughput. Our results may offer some insights into the efficient scheduling of low-latency wireless communications, in which statistical latency and reliability metrics are crucial.
Index Terms:
Low-latency communications, reliability-latency-rate tradeoff, finite-blocklength coding, effective capacity, delay-violation probability
I Introduction
With the tremendous development of communication technologies, the flow of wireless data transmission has experienced an exponential growth in recent years supported by a higher transmission rate and ubiquitous connectivity. Under this trend, multiple delay-sensitive applications have been gaining significant attention in the fifth-generation (5G) communication networks. Besides, the ongoing works for six-generation (6G) communication networks indicate that there will be more stringent delay and reliability requirements in 6G applications [1]. As one of the essential requirements for mission-critical and emerging applications in 5G and 6G, low-latency communication has been a popular topic under active consideration in the research community.
Among the research for low-latency communications, ultra-reliable and low latency communications (URLLC) is a typical application scenario. URLLC aims to achieve an ultra-low and bounded delay to support the above applications. For the next generation URLLC, the requirement on latency will be more stringent [2]. Additionally, to provide real-time experiences for users, communication latency should be controlled or deterministic [3]. However, due to the fading nature of wireless channels, guaranteeing a bounded delay is usually challenging [4], [5]. Besides, limited maximum power brings more difficulty in achieving a bounded delay [6]. Therefore, statistical quality-of-service (QoS) becomes one of the most common choices as the reliability and latency metrics for URLLC [7]. Statistical QoS is characterized by the queue length or delay thresholds and the corresponding queue length or delay violation probability [8]. As mentioned in [9], focusing solely on average-based metrics without considering the tail behavior of wireless systems makes it difficult to satisfy the low-latency and reliability requirements. There are some works focusing on the characterization of the statistical delay in closed-loop control [10] and edge computing systems [11]. Thus, it is crucial to pay attention to statistical QoS analysis in URLLC and other low-latency systems to characterize the performance limits, which can, in turn, help optimize the system’s QoS performance.
To analyze the QoS of communication systems, the first step is to formulate a proper model, which takes QoS metrics into account. In low-latency communications, we are interested in the probability of the queue length and delay violating certain thresholds [9], which are two key performance indicators of QoS. To this end, the effective bandwidth (EB) and effective capacity (EC) methods are proposed to conduct an asymptotic QoS analysis by introducing the link-layer queueing model. With the help of EB and EC, the queue-length-violation probability (QVP) and delay-violation probability (DVP) can be expressed concisely. The EB of a source is defined as the minimum constant service rate required by the given arrival process to satisfy the QoS requirements [12]. Effective capacity (EC) is the dual concept of EB, indicating the maximum constant source rate that can be supported by the given service process to meet the QoS requirement [13]. Based on EB, EC, and Shannon’s formula, several works focus on the performance analysis and resource allocation designs under statistical QoS constraints in various communication systems with different channel and arrival models, such as independent Rayleigh fading channel [14], time-correlated Rayleigh fading channel [15], and Nakagami-$m$ channel [16]. Specifically, the high-SNR slope of EC was characterized in [14] with the infinite blocklength assumption.
With a proper analysis model, the next step is to formulate the appropriate relationship between rate and reliability in the FBL regime. A commonly quoted requirement for 5G is a latency of less than 1 ms and a packet loss probability that is less than or equal to $10^{-5}$ [7]. Given the limited bandwidth, the assumption that the blocklength shall tend to infinity has to be revised to satisfy these requirements. In particular, the traditional Shannon’s formula with an infinite blocklength is not accurate enough in this regime. Thus, finite-blocklength coding (FBC) is expected to play a central role in addressing this problem. A milestone work [17] presented the maximum coding rate in the FBL regime. The normal approximation was also proposed in [17] to approximate the maximum coding rate for the AWGN channel in the FBL regime. This conclusion inspired further studies on the coding rate of various communication scenarios in the FBL regime, such as Gilbert-Elliott channels [18], quasi-static multiple-antenna fading channels [19], and coherent block fading channels [20]. Besides, Lancho, Koch, and Durisi proposed a high-SNR approximation of the maximum coding rate in the FBL regime based on unitary space-time modulation [21]. With the FBL assumption, EC’s properties differ from those in scenarios with the infinite blocklength assumption.
Combining the FBC analysis with the EC-based model, many works tried to characterize and optimize the performance of low-latency communication systems in terms of reliability, latency, and rate. The authors of [22] validated that the DVP derived from LDT is the upper bound of the actual DVP for certain types of arrival processes, even with a small delay threshold. This result indicated the potential use of the EC-based model in URLLC systems. Based on this conclusion, a joint uplink and downlink resource configuration problem was formulated in [23] to improve the spectrum usage efficiency while ensuring reliability and latency performance. In contrast to these works, another line of research focuses on maximizing EC to improve QoS-constrained throughput in the FBL regime. EC maximization has been studied in various fading channels, including Rayleigh [24, 25], Rician [26], and Nakagami-$m$ [27], each with different constraints and conditions in the FBL regime. In [28], the author determined the throughput of a simple ARQ mechanism in the FBL regime, which was applied in [29]. Although many studies have adopted EC with FBC for QoS optimization in low-latency communication systems, the performance limit of EC in the FBL regime remains an open question. Additionally, due to the channel dispersion term in the FBC rate, deriving the analytical expression of EC is not straightforward. This complexity also poses challenges in evaluating the QoS-constrained throughput of low-latency systems and further optimizing EC.
As mentioned above, although many works focus on optimizing the reliability, latency, and service rate in low-latency systems, there is a lack of comprehensive analysis of the fundamental tradeoff, i.e., reliability-latency-rate tradeoff, in the FBL regime. In our previous work [30], the gain-conservation equation was derived to characterize the relationships between the reliability gain, real-time gain, and service-rate gain. However, in [30], only bounded random arrivals were considered, which is equivalent to the deterministic arrival under the high-SNR assumption. Moreover, the conclusions in [30] did not discuss the influence of high SNR on the FBC. Thus, there is still a need for rigorous analysis of the reliability-latency-rate tradeoff in low-latency communication systems with FBC and random arrivals.
In this paper, we characterize the fundamental tradeoff between reliability, latency, and rate in low-latency systems with FBC. Specifically, we start the analysis from the analysis for the high-SNR slope of EC in the FBL regime. We consider two classical scenarios including transmission over an AWGN channel and a Nakagami-$m$ fading channel. First, we consider the case in which the error probability is fixed to determine the service-rate gain, which is the derivative of the EC to the logarithm of the SNR. Based on this result, we further consider the case in which the error probability is a function of SNR. The reliability gain and real-time gain are then defined as the derivatives of logarithms of error probability and DVP to the logarithm of the SNR. Then, the revised gain-conservation equation is characterized over the AWGN channel and fading channel. Moreover, a discussion of the tradeoff between the error probability and latency is provided for the AWGN channel. For efficiently evaluating the EC in the FBL regime, a Laplace’s-method-based EC approximation approach is proposed to derive the analytical expression of EC. Based on the discussion on the fixed-power allocation scheme, we further propose the EC approximation for certain kinds of channel-gain-based power allocation schemes, while its generalization to the systems with the simple retransmission mechanism is also presented.
The rest of this paper is organized as follows. Section II-A presents the system model, including the physical layer model and network layer model. Section II-B introduces the main results of this paper, which are summarized by four main theorems. In Section III A-C, we provide conclusions and corresponding analysis for the revised gain-conservation equations. The tradeoff between the error probability and latency in the AWGN channel is discussed in Section III-D. Further, the EC approximation approach is proposed in Section IV. Numerical results are presented in Section V. Finally, the conclusion is given in Section VI.
Throughout this paper, $u^{\prime}(x)$ denotes the first-order derivative of $u(x)$ for $x$, while $u^{\prime\prime}(x)$ denotes the second-order derivative. $\mathbb{N}^{+}$ denotes the set of positive natural numbers. $\mathbb{E}_{X}\{X\}$ is the expectation of random variable $X$, while ${\rm Var}_{X}\{X\}$ is its variance. For $X[n]$, we omit index $n$ in the subscripts. $e$ denotes Euler’s number. $\boldsymbol{I}_{N}$ denotes a $N$-dimensional identity matrix. For two random variables $X$ and $Y$, ${\rm Cov}(X,Y)$ denote the covariance between $X$ and $Y$. The Gaussian $Q$-function is $Q(x)=\frac{1}{\sqrt{2\pi}}\int_{x}^{\infty}e^{-\frac{t^{2}}{2}}{\rm d}t$, while $Q^{-1}(\cdot)$ represents the inverse Gaussian $Q$-function. $\Gamma(s)=\int_{0}^{\infty}t^{s-1}e^{-t}{\rm d}t$ denote the gamma function. Besides, define the incomplete gamma function as $\Gamma(s,x)=\int_{x}^{\infty}t^{s-1}e^{-t}{\rm d}t$, and the exponential integral function as ${\rm Ei}(x)=-\int_{-x}^{\infty}\frac{e^{-t}}{t}{\rm d}t$.
TABLE I: Main Notation
Symbol Definition
$|h[n]|^{2}$ channel power gain
$T$ the number of time slots in a frame
$f(\cdot)$ distribution of $|h[n]|^{2}$
$m$, $\Omega$ parameters of Nakagami-$m$ fading
$\gamma$ transmitted SNR
$\epsilon$ error probability
$N$ blocklength
$R_{N}^{\epsilon}[n]$ maximum coding rate
$\tilde{R}_{N}^{\epsilon}[n]$ normal approximation
$G(\cdot,\cdot)$ error term of normal approximation
$\theta$ QoS exponent
$\alpha_{A}(\theta)$ effective bandwidth
$\alpha_{S}(\theta)$ effective capacity
$\Lambda(\theta)$ normalized effective capacity
$\chi$ queue-length-violation probability
$\delta$ delay-violation probability
$\Psi(\gamma)$ function of $\gamma$, $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$
$\varrho$ revised QoS exponent
$s_{\infty}$ high-SNR slope of normalize EC
$\zeta$ service-rate gain
$\varpi$ reliability gain
$\tau$ real-time gain
$\Xi(\cdot)$ power allocation scheme
II System Model and Main Results
Figure 1: System Model.
In this section, we present a system model with FBC. Since we mainly focus on the statistical delay performance of this system, we adopt QVP and DVP as the metrics for evaluating the latency performance. More specifically, we will first describe the system model in Section II-A, which includes physical layer and queueing models. Then, we will introduce the main results of this paper, summarized by four main theorems in Section II-B.
II-A System Model
As shown in Fig. 1, we focus on a point-to-point low-latency communication system. The arrival data generated by the source will be transmitted through the channel. Time is divided into time slots with equal length $T_{0}=\frac{N}{B}$ (in seconds), where $B$ is the bandwidth of this system and $N$ is the blocklength. Let $T_{f}$ (in seconds) denote the frame length. Each frame contains $T=\frac{T_{f}}{T_{0}}$ time slots, where $T\in\mathbb{N}^{+}$. [^1] Note that, for the high-mobility or THz communication system, the channel coherence time is significantly small [31]. Thus, $T=1$ also covers the low-latency scenarios. Let $h[n]$ denote the channel coefficient of time slot $n$. Assume that the frame duration is equal to the channel coherence time. Thus, the channel coefficient remains constant within a frame, i.e., $h[iT+1]=h[iT+2]=\cdots=h[(i+1)T]$, $i=0,1,\cdots$, and varies in an independently and identically distributed ($i.i.d.$) manner across different frames. With blocklength $N$, the channel output $\boldsymbol{y}[n]$ at the $n$-th time slot with the input $\boldsymbol{x}[n]$ can be expressed as
$\boldsymbol{y}[n]=h[n]\boldsymbol{x}[n]+\boldsymbol{z}[n],$ (1)
where $\boldsymbol{z}[n]\sim\mathcal{CN}(0,N_{0}\boldsymbol{I}_{N})$ denotes the AWGN. For the AWGN channel, $\boldsymbol{x}[n]\in\mathbb{C}^{N}$ and $h[n]=1$. For the fading channel, the probability density function ($p.d.f.$) of $|h[n]|^{2}$ is denoted by $f(x)$, while its cumulative distribution function ($c.d.f.$) is denoted by $F(x)$. For a wireless system in a Nakagami-$m$ fading channel with $\mathbb{E}_{|h|^{2}}\left\{|h[n]|^{2}\right\}=\Omega$, we have
$f(x)=\left(\frac{m}{\Omega}\right)^{m}\frac{x^{m-1}}{\Gamma(m)}e^{-\frac{m}{\Omega}x},\quad x>0,\,m\geq 0.5.$ (2)
With $\Omega=1$ and $m=1$, the system is in a Rayleigh fading channel with $\mathbb{E}_{|h|^{2}}\left\{|h[n]|^{2}\right\}=1$. For a wireless system with spatial or frequency diversity $\kappa$ in Rayleigh fading channels, the equivalent channel power gain by adopting maximum-ratio combining can be represented by setting $m=\kappa$ and $\Omega=\kappa$.
Next, we introduce the notion of the channel code. In this paper, we assume that the channel state information (CSI) is perfectly known at both the receiver and the transmitter (CSIRT) through the channel estimation and feedback as shown in Fig. 1.222To ensure the perfect CSI assumption, the average power of pilot symbols should be high or the pilot length should be large [32]. With a large coherence bandwidth or coherence time, the pilot length can be large, such as in indoor scenarios with small delay and Doppler spreads [33, 34]. Additionally, to achieve perfect CSIT, there must be a feedback path with a high-resolution quantization scheme and high SNR to transmit CSI precisely and reliably [34]. For latency-constrained systems, available time resources for transmission is limited. Thus, using high SNR or considering systems with a large coherence bandwidth can ensure the perfect CSI assumption while meeting the latency constraints. $\Delta N=B\Delta T$ channel uses are utilized for the channel estimation in each frame, which we neglect in this paper. [^3] Note that for short packet communication with extremely small blocklength, the assumption that perfect CSIRT may be unrealistic. Moreover, analyzing the system with extremely small blocklength requires more effort to derive a precise expression for the coding rate based on [17]. With CSIRT, the transmitter can adaptively determine the coding rate. Specifically, with $h[n]=h$, an $(M_{h},N,\gamma,\epsilon)$ code consists of [35]
  1. 1.
    An encoder $f^{h}_{\rm tx}:\{1,\cdots,M_{h}\}\times h\to\mathbb{C}^{N}$ that maps the message $j\in\{1,\cdots,M_{h}\}$ and the channel $h$ to a codeword $\boldsymbol{x}=f^{h}_{\rm tx}(j,h)$ satisfying
    $\frac{\|\boldsymbol{x}[n]\|^{2}}{N_{0}}=N\gamma.$ (3)
  2. 2.
    A decoder $g_{\rm rx}^{h}:\mathbb{C}^{N}\times h\to\{1,\cdots,M_{h}\}$ satisfying
    ${\color[rgb]{0,0,0}\Pr\left\{g_{\rm rx}^{h}(\boldsymbol{x},h)\neq J|J\in\{1,\cdots,M_{h}\}\right\}\leq\epsilon.}$ (4)
The maximum instantaneous coding rate in time slot $n$, which is defined as $R_{N}^{\epsilon}[n]$, with $(M_{h},N,\gamma,\epsilon)$ code is thus defined as
$R_{N}^{\epsilon}[n]=\sup\left\{\log_{2}M_{h}:\forall(M_{h},N,\gamma,\epsilon)\text{ code}\,|\,h[n]=h\right\}.$ (5)
The conclusion in [17, 36] showed that for channels with capacity $C$, the instantaneous maximum coding rate (in bits per $N$ channel uses) with $\gamma$, $N$, and $\epsilon$, can be expressed as
$R_{N}^{\epsilon}[n]=NC[n]-\sqrt{NV[n]}Q^{-1}(\epsilon)+\frac{\log N}{2}+O\left(1\right),$ (6)
where $V[n]$ is the channel dispersion, and $O(1)$ is the error term that is bounded for large $N$ and may be related to $\gamma$ and $\epsilon$ at the receiver [17]. [^4] Note that there is no existing work presenting a general conclusion on whether the error term is bounded for $\gamma$. We will discuss this in Section II-B. For the AWGN channel, we omit the index $n$. According to [36], the maximum coding rate (in bits per $N$ channel uses) for a given $\epsilon$, $N$, and $\gamma$ for the AWGN channel is given by
${\color[rgb]{0,0,0}R_{N}^{\epsilon}=N\log_{2}(1+\gamma)-\sqrt{N\left(1-\frac{1}{(1+\gamma)^{2}}\right)}Q^{-1}(\epsilon)\log_{2}e+\frac{\log_{2}N}{2}+G(N,\gamma,\epsilon),}$ (7)
where we call $G(N,\gamma,\epsilon)$ error term function. For the given $\gamma$ and $\epsilon$, $G(N,\gamma,\epsilon)$ satisfies [36]
${\color[rgb]{0,0,0}\lim_{N\to+\infty}|G(N,\gamma,\epsilon)|<+\infty.}$ (8)
Moreover, given the channel gain $|h[n]|^{2}$, the maximum coding rate (in bits per $N$ channel uses) with $\epsilon$, $N$, and $\gamma$ for the fading channel is given by
${\color[rgb]{0,0,0}R^{{}_{\epsilon}}_{N}[n]\!\!=\!\!N\log_{2}\!\left(1\!+\!|h[n]|^{2}\gamma\right)\!-\!\sqrt{\!N\!\left(1\!-\!\frac{1}{\left(1+|h[n]|^{2}\gamma\right)^{2}}\right)}Q^{-1}(\epsilon)\log_{2}e+\frac{\log_{2}N}{2}+G\left(N,|h[n]|^{2}\gamma,\epsilon\right).}$ (9)
For simplicity, we define $\tilde{R}^{\epsilon}_{N}[n]=R^{\epsilon}_{N}[n]-G(N,|h[n]|^{2}\gamma,\epsilon)$, which is called normal approximation in [17].
To backlog the packets to be transmitted, an infinite-length buffer is assumed to be available at the transmitter. Let $a[n]$ denote the size of the arrival data at the beginning of time slot $n$. In this paper, we assume that $a[n]$ follows an arbitrary stochastic distribution with finite expectation and variance in this system, and $a[n]$ changes in an $i.i.d.$ manner across different time slots. The average size of the arrival data is denoted by $\mu_{a}$, while the variance is denoted by $\sigma_{a}^{2}$. Accordingly, the accumulated number of the arrival data until time slot $M$ can be expressed as $A[M]=\sum_{n=1}^{M}a[n]$. Similarly, $s[n]$ is defined as the size of the service data in time slot $n$. The average size of the service data is denoted by $\mu_{s}$, while the variance of $s[n]$ is denoted by $\sigma_{s}^{2}$. The accumulated number of service data until time slot $M$ can be expressed as $S[M]=\sum_{n=1}^{M}s[n]$.
As shown in Fig. 1, we assume a first-in first-out (FIFO) policy for serving the backlogged packets. Let $Q[n]$ denote the length of the queue at the end of time slot $n$. Then, the queue length is updated as
$Q[n]=\left(Q[n-1]+a[n]-s[n]\right)^{+},$ (10)
where $(x)^{+}=\max\{0,x\}$.
Remark 1: In this paper, we mainly focus on the analysis with CSIRT. There are also some conclusions on the coding rate in the FBL regime with CSIR only. The rigorous analysis of $R_{N}^{\epsilon}$ in the block fading channel with CSIR for single-input and single-out (SISO) systems was performed in [37]. With real transmitted symbols and noise, an analytical expression of the dispersion was given in Eqs. (27) and (36) of [37] under the assumption that each symbol experiences $i.i.d.$ fading. For simplicity, we omit the index $n$ in this remark. The dispersion for the real block fading channel is given by
${\color[rgb]{0,0,0}V_{r}={\rm Var}_{|h|^{2}}\left(\frac{1}{2}\log_{2}\left(1+|h|^{2}\gamma\right)\right)+\frac{\log_{2}e}{2}\left(1-\mathbb{E}_{|h|^{2}}^{2}\left\{\frac{1}{1+|h|^{2}\gamma}\right\}\right).}$ (11)
Moreover, the dispersion for the complex block fading channel is given by
${\color[rgb]{0,0,0}V_{c}={\rm Var}_{|h|^{2}}\left(\log_{2}\left(1+|h|^{2}\gamma\right)\right)+\log_{2}e\left(1-\mathbb{E}_{|h|^{2}}^{2}\left\{\frac{1}{1+|h|^{2}\gamma}\right\}\right).}$ (12)
This conclusion is generalized in [20], which considered multiple-input and multiple-output (MIMO) systems with real transmitted symbols and noise. Additionally, they considered the case in which the $T_{d}$ symbols experience the same fading. According to Eqs. (16) and (17) in [20], the dispersion $V$ for a SISO system with a real block fading channel is denoted by
$V_{m}=$ $T_{d}{\rm Var}_{|h|^{2}}\left(\frac{1}{2}\log_{2}\left(1+|h|^{2}\gamma\right)\right)+\frac{\log_{2}^{2}e}{2}\mathbb{E}_{|h|^{2}}\left\{1-\frac{1}{(1+|h|^{2}\gamma)^{2}}\right\}+\gamma^{2}\left(\eta_{1}-\frac{\eta_{2}}{T_{d}}v^{*}(1,T_{d})\right)$ (13)
$=$ $T_{d}{\rm Var}_{|h|^{2}}\left(\frac{1}{2}\log_{2}\left(1+|h|^{2}\gamma\right)\right)+\frac{\log_{2}^{2}e}{2}\left(1-\mathbb{E}_{|h|^{2}}^{2}\left\{\frac{1}{1+|h|^{2}\gamma}\right\}\right),$
where $\eta_{1}$, $\eta_{2}$, $v^{*}(1,T_{d})$ are provided in Eqs. (14), (15), and (23) of [20], respectively.
By comparing the results from [37] and [20], we can see that the result in [20] is a generalization of [37] for the real block fading channel. Given $T_{d}=1$, they are consistent. For more general cases, e.g., $T_{d}>1$ in SISO systems or MIMO systems with real block fading channel, the result in [20] provides further insights.
II-B Effective Bandwidth and Capacity
In this subsection, we will introduce effective bandwidth (EB) and effective capacity (EC). The essence of EB and EC is to use LDT to characterize the tail distribution of the length and delay violation probability. By using the Gartner-Ellis theorem [38] and given a length threshold $L$, QVP satisfies
$\lim_{L\to+\infty}\frac{\ln\Big{(}\sup_{m}\big{\{}\Pr\{Q[m]\geq L\}\big{\}}\Big{)}}{L}=-\theta,$ (14)
where $\theta$ is called QoS exponent, indicating the decay rate of the tail distribution of the queue length. According to [12], the EB of the arrival process is defined as
${\color[rgb]{0,0,0}\alpha_{A}(\theta)=\lim_{M\to+\infty}\frac{1}{\theta M}\ln\mathbb{E}_{A}\left\{e^{\theta A[M]}\right\}.}$ (15)
The EB of the arrival process indicates the minimum constant transmission rate for guaranteeing certain QoS requirements as shown in Eq. (14). Similarly, according to [13], the EC of the service process is defined as
${\color[rgb]{0,0,0}\alpha_{S}(\theta)=\lim_{M\to+\infty}-\frac{1}{\theta M}\ln\mathbb{E}_{S}\left\{e^{-\theta S[M]}\right\}.}$ (16)
The EC of the service process refers to the maximum constant arrival rate that meets the QoS requirements shown in Eq. (14).
With a given $\gamma$, $s[n]$ is $i.i.d.$. Thus, for $T\geq 1$, EC is given by
$\alpha_{S}(\theta)=$ $\lim_{M\to+\infty}-\frac{1}{\theta M}\ln\mathbb{E}_{S}\left\{e^{-\theta S[M]}\right\}$ (17)
$=$ $\lim_{M\to+\infty}-\frac{1}{\theta T}\cdot\frac{T}{M}\ln\mathbb{E}_{s}\left\{e^{-\theta\sum_{i=0}^{\frac{M}{T}-1}\sum_{j=1}^{T}s[iT+j]}\right\}$
$=$ $-\frac{1}{\theta T}\ln\mathbb{E}_{|h|^{2}}\left\{e^{-\theta Ts[n]}\right\}.$
To simplify the derivation, we normalize the EC by the blocklength, which is also called effective rate [14]. We let $\Lambda(\gamma)$ denote the normalized EC, which is a function of $\gamma$ with a given $\theta$. According to Eq. (17), $\Lambda(\gamma)$ of the $i.i.d.$ $s[n]$ is given by
[…]
Note that the effective rate defined in [14] is equal to $\Lambda(\gamma)$ when $T=1$. In this paper, we use $\Lambda(\gamma)$ to characterize the service-rate gain of the low-latency communication systems.
Let $Q[\infty]$ denote the steady state of $Q[n]$. If $Q[\infty]$ exists, QVP is denoted by $\chi=\Pr\{Q[\infty]\geq L\}$. With a large threshold $L$ (in bits), $\chi$ can be approximated according to Eq. (14) and [13] as
$\chi\simeq\eta e^{-\theta L}.$ (19)
In Eq. (19), $\eta$ is the probability that the queue is not empty, which can be approximated by the ratio of $\mu_{a}$ to $\mu_{s}$. The notation $a(x)\simeq b(x)$ means that $\lim_{x\to\infty}\frac{a(x)}{b(x)}=1$. In this paper, we focus on the high-load scenario, i.e., $\eta\to 1$, which is more common in emerging applications with the increasing demand for information exchange [39] and is more likely to cause the occurrence of extreme events, e.g., queue length and delay violations  [^5] Note that by setting $\eta=1$ we obtain upper bounds of $\chi$ and $\delta$ defined below with arbitrary loads [40]. Thus, the derived results in this paper can be seen as a converse analysis for the low-load scenario.. $\theta$ is the solution to the following equation [41]:
$\alpha_{A}(\theta)-\alpha_{S}(\theta)=0.$ (20)
Let $D[m]$ denote the delay experienced by the data arriving at the start of time slot $m$. The steady state of $D[n]$ is denoted by $D[\infty]$. If $D[\infty]$ exists, DVP is denoted by $\delta=\Pr\{D[\infty]\geq D_{\rm max}\}$. With a large delay bound $D_{\rm max}$ (in slots), $\delta$ in the high load scenario is approximated as [13], [40]
$\delta\simeq e^{-\theta\alpha_{A}(\theta)D_{\rm max}}.$ (21)
II-C Main Results
Figure 2: Structures of Section III and Section IV.
In this subsection, we present the main results of this paper. The main results of this paper can be divided into two aspects, which refer to the revised gain-conservation equation with random arrivals in the high-SNR regime and a Laplace’s-method-based EC approximation approach. The detailed analysis and discussion are provided in Sections III and IV. The structure of Sections III and IV are shown in Fig. 2.
Before presenting the main results, we first give some definitions. We define $\zeta=\lim\limits_{\gamma\to+\infty}\frac{\partial{\Lambda(\gamma)}}{\partial{\log_{2}\gamma}}$ as the QoS-constrained service-rate gain, similar to the definition of the service-rate gain in [30]. $\zeta$ indicates the rate of increase of throughput with the increase of $\log_{2}\gamma$ in the high-SNR regime. Let $\varpi=\sqrt{2\log_{2}e}\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\sqrt{N}\partial\log_{2}\gamma}$ denote the reliability gain. $\varpi$ is related to the decay rate of error probability with the increase of $\log_{2}\gamma$ in the high-SNR regime, consistent with the reliability gain in [30]. [^6] Note that the constant $\sqrt{2\log_{2}e}$ is also consistent with [30] since dispersion is treated as a constant in [30]. We can also define $\varpi$ without this constant, which will not affect the main parts of the proposed conclusions. The only difference is that the coefficient of $\varpi$ in the gain-conservation equation will change. We also define $\tau=-\frac{1}{D_{\rm max}}\frac{\partial\log_{2}\delta}{\partial\log_{2}\gamma}$ as the real-time gain, consistent with the definition in [30]. $\tau$ represents the decay rate of DVP with the increase of $\log_{2}\gamma$ in the high-SNR regime. We refer to $\tau$ as real-time gain because real-time usually refers to being controlled and deterministic [3]. $\tau$ represents the decay rate of the probability of extreme events happening. However, $\tau$ can also be named differently, e.g., revised diversity gain  [^7] Diversity gain was originally proposed in [42]. We will present the comparison between $\tau$ and diversity gain in Section III-C., low-latency gain, or timeliness gain. Based on the above definitions, the high-SNR reliability-latency-rate tradeoff is shown in Theorem 1 for low-latency communication systems with FBC and random arrivals. Let $\varrho$ denote an arbitrarily positive constant. The main results and quantities related to gain-conservation equations are summarized in Fig. 3.
Figure 3: Main results for gain-conservation equations and relationships between different quantities.
Theorem 1. Assume $\epsilon\in(0,0.5]$ is a function of $\gamma$ and $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$. Under the condition that $\varpi$ is finite, $\theta N=\varrho$, and $N=\Psi(\gamma)$ satisfies
${\color[rgb]{0,0,0}\sup_{\gamma>0}\left|\frac{G\left(\Psi(\gamma),|h[n]|^{2}\gamma,\epsilon(\gamma)\right)}{\Psi(\gamma)}\right|\leq\nu,\quad\text{for all}|h[n]|^{2},}$ (22)
where $\nu\geq 0$ is a finite constant,
  • for a SISO system in an AWGN channel, we have
    $\zeta+c\varpi=1,\,\text{and}$ (23)
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (24)
    By combing Eqs. (23) and (24), we have
    $(1-\theta N\log_{2}e)\zeta+c\varpi+\tau=1;$ (25)
  • for a SISO system in a Nakagami-$m$ fading channel with a finite $\varpi$, we have
    $\zeta+c\varpi=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\},\,\text{and}$ (26)
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (27)
    By combing Eqs. (26) and (27), we have
    $(1-\theta N\log_{2}e)\zeta+c\varpi+\tau=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\}.$ (28)
In Eqs. (25) and (28), $c\in[0,1]$ is given by
$c=\lim_{\gamma\to+\infty}\frac{1}{1+\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}}.$ (29)
Proof:
This theorem summarizes the main results in Section III. The proof of this theorem follows from the proofs of Theorem 5, Theorem 6, Lemma 1, and Theorem 7 (in Appendix D, Appendix E, Appendix F, and Section III-C, respectively).
From the definitions of $\zeta$, $\varpi$, and $\tau$, we find that these three gains characterize the communication system’s performance in terms of the service rate, reliability, and latency with a maximum power constraint. Thus, this theorem presents a fundamental tradeoff among the service rate, reliability, and latency in the high-SNR regime for low-latency communications. We combine Eqs. (26) and (27) to derive (28) so that we can compare Eq. (28) with the Eq. (4) of [30], which has a similar form. In this work, we adopt EC to characterize the service-rate and real-time gains, which is different from the asymptotic analysis in [30]. The main difference between Eqs. (25), (28), and Eq. (4) in [30] results from the influence of random arrivals. Thus, Theorem 1 can be seen as a generalization of Eq. (4) in [30] to the systems with unbounded random arrivals under a large delay threshold. Moreover, [30] lacks a discussion of the assumptions underlying the use of normal approximation in asymptotic analysis, which are provided in this paper.
The assumption shown in Eq. (22) can be met with $G$ function. For simplicity, we omit the index $n$ in the following analysis. Besides, we let $V\left(|h|^{2}\gamma\right)=\left(\log_{2}e\right)^{2}\left(1-\left(1+|h|^{2}\gamma\right)^{-2}\right)$ denote the dispersion term. In [17], Polyanskiy, Poor, and Verdú proposed the upper and lower bounds of $R_{N}^{\epsilon}$, which are given by  [^8] There is a tighter lower bound proposed in [36], which shows that $NC-\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(\epsilon)+\frac{1}{2}\log_{2}N+O(1)\leq R_{N}^{\epsilon}$. We do not use this bound since the $O(1)$ term in this bound is difficult to analyze. Besides, the lower bound proposed in [18] is sufficient for our analysis.
${\color[rgb]{0,0,0}NC-\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(\epsilon)+O(1)\leq R_{N}^{\epsilon}\leq NC-\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(\epsilon)+\frac{\log_{2}N}{2}+O(1),}$
where the upper bound requires $N>\left(\frac{2K\left(|h|^{2}\gamma\right)}{(1-\epsilon)V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)^{2}$ and the lower bound requires $N>\left(\frac{2K\left(|h|^{2}\gamma\right)}{\epsilon V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)^{2}$. Based on the upper and lower bounds of $R_{N}^{\epsilon}$, we can obtain the upper and lower bounds of $G$ function, which are denoted by $g_{u}(|h|^{2}\gamma,\epsilon)$ and $g_{l}(N,|h|^{2}\gamma,\epsilon)$, respectively. Specifically, according to Eq. (632) in [17], the upper bound of $R_{N}^{\epsilon}$ is given by
$R_{N}^{\epsilon}$ $\leq NC+\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(1-\epsilon)+\frac{1}{2}\log_{2}N+g_{u}\left(|h|^{2}\gamma,\epsilon\right)$ (30)
$=\tilde{R}_{N}^{\epsilon}+\frac{c_{1}}{V\left(|h|^{2}\gamma\right)}+\frac{3}{2}\log_{2}V\left(|h|^{2}\gamma\right)-\log_{2}K\left(|h|^{2}\gamma\right).$
In Eq. (30), $K(\cdot)$ is defined as
${\color[rgb]{0,0,0}K\left(|h|^{2}\gamma\right)=c_{0}\left(\frac{|h|^{2}\gamma}{1+|h|^{2}\gamma}\right)^{3}\mathbb{E}_{z}\left\{\left|z^{2}-\frac{2}{\sqrt{|h|^{2}\gamma}}z-1\right|^{3}\right\},}$ (31)
where $c_{0}$ is a positive constant and $z$ is the standard normal variable. $c_{1}$ in Eq. (30) is given by
${\color[rgb]{0,0,0}c_{1}=-2K\left(|h|^{2}\gamma\right)\cdot\min_{y\in\left[1-\epsilon-b_{1},1-\epsilon\right]}\frac{{\rm d}Q^{-1}(y)}{{\rm d}y},}$ (32)
where $b_{1}$ is given by
${\color[rgb]{0,0,0}b_{1}=\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)\sqrt{\left(\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)(1-\epsilon)}\right)^{2}+1}}.}$ (33)
According to Eqs. (650)-(654) in [17], the lower bound of $R_{N}^{\epsilon}$ is given by
$R_{N}^{\epsilon}$ $\geq NC+\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(1-\epsilon)+\frac{1}{2}\log_{2}N+{\color[rgb]{0,0,0}g_{l}(N,\gamma,\epsilon)}$ (34)
${\color[rgb]{0,0,0}=\tilde{R}_{N}^{\epsilon}-\frac{1}{2}\log_{2}N+\log_{2}\frac{K\left(|h|^{2}\gamma\right)}{c_{2}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}-\log_{2}\left[2\left(\frac{\ln 2}{\sqrt{2\pi}}+\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)\right]}$
${\color[rgb]{0,0,0}\quad\quad\quad\quad\quad\quad+\sqrt{NV\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{N}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon)\right],}$
where $c_{2}>0$ is a constant. Based on Eqs. (30) and (34), we can obtain the upper and lower bounds of $G(N,|h|^{2}\gamma,\epsilon)$, respectively. Thus, the requirements on $\Psi(\gamma)$ to satisfy Eq. (22) can be determined with the help of Eq. (22) through Eqs. (30) and (34). In Theorems 2 and 3, we present the sufficient conditions with which Eq. (22) holds under different system assumptions.
Theorem 2. Given $\epsilon\in(0,0.5]$, the condition proposed in Eq. (22) is satisfied if $\Psi(\gamma)$ follows
${\color[rgb]{0,0,0}\Psi(\gamma)>\varsigma_{1},\quad\gamma>0,}$ (35)
where $\Psi(\gamma)$ can be bounded or satisfy $\lim_{\gamma\to+\infty}\Psi(\gamma)=+\infty$. $\varsigma_{1}$ is finite and is given by
${\color[rgb]{0,0,0}\varsigma_{1}=\max_{x\geq 0}\left\{\left(\frac{2K(x)}{(1-\epsilon)V^{\frac{3}{2}}(x)}\right)^{2},\left(\frac{2K(x)}{\epsilon V^{\frac{3}{2}}(x)}\right)^{2}\right\}.}$ (36)
Proof:
See Appendix A.
From Theorem 2 we find that with a given $\epsilon$, $\Psi(\gamma)$ can be set as a constant larger than $\varsigma_{1}$. Moreover, it can also be chosen as $\Psi(\gamma)=\varsigma_{1}+U(\gamma)$, where $U(x)>0$ for $x>0$ and $\lim\limits_{x\to+\infty}U(x)=+\infty$. Theorem 2 is used to determine the forms of $\Psi(\gamma)$ in Theorem 5 and Corollary 1 in Section III-A.
In the proof of Theorem 2, we have also found that, given $N>\varsigma_{1}$ and $\epsilon$, the error term $G(N,\gamma,\epsilon)$ is bounded for $\gamma$, which supplements to [21]. An important future work is to further discuss the case with $N<\varsigma_{1}$.
Theorem 3. With the assumption that $\epsilon\in(0,0.5]$ is a function of $\gamma$ and $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$, the condition proposed in Eq. (22) is satisfied if $\Psi(\gamma)$ follows
$\begin{cases}&\Psi(\gamma)\epsilon^{2}(\gamma)>\varsigma_{2},\quad\gamma>0,\\ &\lim_{\gamma\to+\infty}\frac{e^{\left[\frac{Q^{-1}(\epsilon(\gamma))}{2}\right]^{2}}}{\Psi(\gamma)}<+\infty,\\ &\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{\Psi(\gamma)}}<+\infty,\end{cases}$ (37)
where $\varsigma_{2}$ is a finite and positive constant, satisfying
${\color[rgb]{0,0,0}\varsigma_{2}=4\max_{x\geq 0}\left\{\frac{K^{2}(x)}{V^{3}(x)}\right\}.}$ (38)
Proof:
See Appendix B.
Theorem 3 presents the specific requirements on $\Psi(\gamma)$ for the cases with $\epsilon(\gamma)\in(0,0.5]$ and $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$. The form of $\Psi(\gamma)$ is related to $\epsilon(\gamma)$. Given $\epsilon(\gamma)$, $\Psi(\gamma)$ can be conceived according to Eq. (37). For example, let $\Psi(\gamma)=e^{\left[\frac{Q^{-1}(\epsilon(\gamma))}{2}\right]^{2}}+\frac{\varsigma_{2}+1}{\epsilon^{2}(\gamma)}$, which satisfies Eq. (37). Theorem 3 is used to determine the forms of $\Psi(\gamma)$ in Theorems 1, 6, 7 and Corollary 2.
In this paper, we define the constant $\varrho$ as the revised QoS exponent. Different from the definition of $\theta$ in Eq. (14), $\varrho$ is related to the value of the DVP. We can better understand the interpretation of $\varrho$ by providing the expression of the DVP. By combining Eqs. (18), (20), and (21), we obtain
${\color[rgb]{0,0,0}\delta\simeq\exp\left(\frac{D_{\rm max}}{T}\ln\mathbb{E}_{|h|^{2}}\left\{\exp\left(-\varrho T\frac{R_{N}^{\epsilon}[n]}{N}\right)\right\}\right).}$ (39)
From Eq. (39), we find that a larger $\varrho$ results in a smaller $\delta$ given $\gamma$, $\Psi(\cdot)$, and $D_{\rm max}$. Thus, the value of $\varrho$ represents the QoS requirements on DVP.
EC is widely used in research on low-latency communications to evaluate the QoS-constrained throughput of wireless systems. However, it is usually challenging to obtain the closed-form expression of EC, which brings difficulty for the designs of resource allocation and scheduling for maximizing EC. In this paper, we propose a Laplace’s-method-based EC approximation approach, in which normal approximation is used to evaluate the maximum coding rate. Thus, this approximation can be used when normal approximation has satisfying accuracy. For simplicity of expression, let
$r(x)=\log_{2}\left(1+\Xi(x)x\right)-\sqrt{\frac{1}{N}\left(1-(1+\Xi(x)x)^{-2}\right)}Q^{-1}(\epsilon)\log_{2}e+\frac{\log_{2}N}{2N},$ (40)
where $x$ denotes the channel power gain $|h[n]|^{2}$ in time slot $n$. Since $|h[n]|^{2}$ is $i.i.d.$ across time slots, we omit the argument for $x$. $\Xi(x)$ denotes the channel-power-gain-based power allocation scheme. Note that $\Xi(x)$ can also denote the fixed power allocation by setting $\Xi(x)=\gamma$. Besides, let $\ell=\left(\frac{Q^{-1}(\epsilon)}{\sqrt{N}}\right)^{2}$.
Theorem 4. For a low-latency communication system that adopts a power allocation scheme $\Xi(x)$ satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ with channel distribution $f(\cdot)$, EC is approximated as
$\alpha_{S}(\theta)=Nr(x^{*})-\frac{\ln f(x^{*})+\frac{1}{2}\ln 2\pi-\frac{1}{2}\ln|r^{\prime\prime}(x^{*})|-\frac{1}{2}\ln\theta NT}{\theta T}+O\left(\frac{1}{\theta^{2}T^{2}N}\right),$ (41)
where $x^{*}$ is given by
$\Xi(x^{*})x^{*}=\sqrt{\frac{1+\sqrt{1+4\ell}}{2}}-1.$ (42)
For fixed power allocation $\Xi(x)=\gamma$, $x^{*}$ is given by
$x^{*}=\frac{1}{\gamma}\left(\sqrt{\frac{1+\sqrt{1+4\ell}}{2}}-1\right).$ (43)
Proof:
Theorem 4 summarizes the main results in Section IV. The proof of this theorem follows from the proof of Lemma 2 and Theorem 8 (in Appendix H and I, respectively).
The approximation proposed in Theorem 4 gives us a way to obtain the closed-form expression of the EC, which can help us further analyze the QoS-constrained performance of low-latency communication systems and compare different scheduling schemes. The condition mentioned in Theorem 4 can be satisfied by some classical allocation schemes. More detailed discussion on $\Xi(x)$ is presented in Section IV-B.
III Reliability-latency-rate tradeoff in the finite blocklength regime
In this section, we mainly focus on characterizing the reliability-latency-rate tradeoff in the FBL regime for low-latency communication systems. In Section III-A, we start by deriving the high-SNR slope of EC in the FBL regime with fixed error probability, which indicates the service-rate gain of the low-latency communication systems. Based on the results of the high-SNR slope of EC in the FBL regime, we further generalize the conclusions to characterize the reliability gain by representing error probability as a function of SNR in Section III-B. Finally, the gain-conservation equation, which reveals the tradeoff between the service-rate gain, reliability gain, and real-time gain in the FBL and high-SNR regime, is proposed in Section III-C.
III-A High-SNR Slope of EC in the FBL regime
In this subsection, we focus on characterizing the high-SNR slope of EC in the FBL regime. Specifically, we derive the high-SNR slope of EC with a given $\epsilon$ in the AWGN and Nakagami-$m$ fading channel with finite blocklength coding. The derived conclusions can be seen as a generalization of conclusions in [14]. We then provide the corollaries for systems in Rayleigh fading channels with and without frequency or spatial diversity.
We start our discussion with the FBL-SISO system in an AWGN channel. Specifically, the high-SNR slope of EC in the FBL regime is defined as [14]
$s_{\infty}=\lim_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}.$ (44)
We then provide the discussion for the FBL-SISO system in the AWGN and Nakagami-$m$ fading channel, which covers wired communications, satellite communications, and common scenarios in wireless communications. In Theorem 5, we derive the high-SNR slope of the normalized EC for the FBL-SISO system in the AWGN and Nakagami-$m$ fading channel.
Theorem 5. Under the condition that $N=\Psi(\gamma)$ satisfies Theorem 2 with a given $\epsilon$,
  • the high-SNR slope of the normalized EC for the FBL-SISO system in the AWGN channel is $s_{\infty}=1$;
  • the high-SNR slope of the normalized EC for the FBL-SISO system in the Nakagami-$m$ fading channel with $m\in[0.5,+\infty)$ is given by
    $\zeta=\begin{cases}1,\quad&1\leq\frac{m}{\theta NT\log_{2}e},\\ \frac{m}{\theta NT\log_{2}e},\quad&1>\frac{m}{\theta NT\log_{2}e}.\end{cases}$ (45)
Proof:
See Appendix D.
In Theorem 5, we obtain the value of $s_{\infty}$ in the FBL-SISO system with the AWGN and Nakagami-$m$ fading channels. We can compare this result with the well-known multiplexing gain [42]. Multiplexing gain is defined as the ratio of the maximum coding rate to $\log_{2}\gamma$ as $\gamma\to+\infty$, which is equal to 1 for the SISO system. The definition of multiplexing gain does not include the influence of latency requirements. As we mentioned in Section II-B, $\varrho=\theta N$ indicates the constraint of DVP, which is one of the key performance indicators of QoS. A larger $\theta N$ indicates a stricter QoS constraint according to Eq. (39). Thus, $s_{\infty}$ can be seen as the revised multiplexing gain, which takes the requirements of DVP into account. Since $s_{\infty}$ considers the gain under the QoS constraint, it should not be larger than the multiplexing gain, as validated in Theorem 5.
The results in Theorem 5 show that with a small $\theta N$, $s_{\infty}$ of the fading channel is independent of the statistical constraints, which is consistent with the analysis in [14] for the infinite-blocklength case in the Rayleigh fading channel. This result indicates that if the requirement of DVP is loose, it does not influence the rate of increase of the QoS-constrained throughput. Furthermore, we find that $s_{\infty}$ of the fading channel can be smaller than 1, which is also consistent with the common understanding that the randomness of the fading channel has a negative influence on the QoS-constrained throughput of queueing systems.
Based on Theorem 5, we can obtain $s_{\infty}$ for many common cases. In Corollary 1, we present the $s_{\infty}$ of Rayleigh fading channels with and without frequency or spatial diversity.
Corollary 1. With the same assumption used in Theorem 5,
  • the high-SNR slope of the normalized EC in the Rayleigh fading channel is given by
    $s_{\infty}=\begin{cases}1,\quad&1\leq\frac{1}{\theta NT\log_{2}e},\\ \frac{1}{\theta NT\log_{2}e},\quad&1>\frac{1}{\theta NT\log_{2}e}.\end{cases}$ (46)
  • the high-SNR slope of the normalized EC in Rayleigh fading channels with frequency or spatial diversity $\kappa$ by adopting maximum-ratio combining is given by
    $s_{\infty}=\begin{cases}1,\quad&1\leq\frac{\kappa}{\theta NT\log_{2}e},\\ \frac{\kappa}{\theta NT\log_{2}e},\quad&1>\frac{\kappa}{\theta NT\log_{2}e}.\end{cases}$ (47)
Proof:
The results can be obtained directly by setting $m=\Omega=1$ and $m=\Omega=\kappa$ from Eq. (45). ∎
In Theorem 5 and Corollary 1, we characterize the high-SNR slope of the normalized effective capacity in the FBL regime with a given $\epsilon$. By comparing the results in Theorem 5 and Corollary 1, we find that in the FBL regime, $s_{\infty}$ of the systems with the frequency or spatial diversity is not lower than that of the SISO system in a Rayleigh fading channel. For the SISO system in a Nakagami-$m$ fading channel, its $s_{\infty}$ is not lower than that of the SISO system in a Rayleigh fading channel with $m\geq 1$, while its $s_{\infty}$ is not larger than that of the SISO system in a Rayleigh fading channel with $m\in[0.5,1)$. Furthermore, we also find that, with a larger $\kappa$ or $m$, the length of the interval in which $s_{\infty}$ is irrelevant to $\theta N$ is larger. According to [33], $m$ measures the ratio of line-of-sight signal power to multipath power, which means $m<1$ corresponds to a worse fading than the Rayleigh fading and $m>1$ corresponds to a less severe fading than the Rayleigh fading. Thus, from the above results, we find that the severity of the fading and the spatial or frequency diversity influence $s_{\infty}$. Less serve fading and larger spatial or frequency diversity help us obtain a larger $s_{\infty}$ with a given QoS constraint and ensure $s_{\infty}$ irrelevant to QoS constraints in a larger $\theta N$ set.
Remark 2: We can also define the $s_{0}$ as the low-SNR slope of the normalized EC in the FBL regime. However, we have to determine the denominator in the definition of $s_{0}$. As far as we know, $\gamma$ is not a proper choice since it can result in $\frac{\Lambda(\gamma)}{\gamma}\to+\infty$ when $\gamma$ approaches 0 by adopting a kind of orthogonal code proposed in Chapter 8.5 of [43]. After determining the proper choice of the denominator, we can obtain the upper bound of $s_{0}$ by considering the infinite blocklength with the help of Shannon’s formula. Performing the above analysis is not trivial, which makes it an important future work.
III-B Reliability Gain in the FBL regime
In this subsection, we further consider the case in which $\epsilon$ is a function of $\gamma$. Based on this assumption, we will refine the conclusions in Theorem 5. The revised results are summarized in Theorem 6.
Theorem 6. Assume $\epsilon\in(0,0.5]$ is a function of $\gamma$ and $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$. Under the assumption that $N=\Psi(\gamma)$ satisfies Theorem 3,
  • for an FBL-SISO system in an AWGN channel, we have
    $\lim_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}+\log_{2}e\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}=1.$ (48)
  • for an FBL-SISO system in a Nakagami-$m$ fading channel, we have
    $\lim_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}+\log_{2}e\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\},$ (49)
Proof:
See Appendix E.
In Theorem 6, we proposed the result of the high-SNR asymptotic analysis of the EC, which is related to the error probability. Before characterizing the tradeoff between the service-rate gain, reliability gain, and real-time gain, we need to further discuss the reliability gain to build the link between them. First, we define the reliability gain. Let $\varpi=\sqrt{2\log_{2}e}\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\sqrt{N}\partial\log_{2}\gamma}$ denote the reliability gain. $\varpi$ is related to the decay rate of error probability with the increase of $\log_{2}\gamma$ in the high-SNR regime, which is consistent with [30]. With this definition, we can derive the relationship between reliability gain and the second term on the left-hand side of Eqs. (48) and (49). We summarize this relationship in Lemma 1.
Lemma 1. Given a finite reliability gain $\varpi$, it satisfies
$\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}=\frac{c_{4}}{\log_{2}e}\varpi,$ (50)
where $c_{4}\in[0,1]$ is given by
$c_{4}=\lim_{\gamma\to+\infty}\frac{1}{1+\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}}.$ (51)
Proof:
See Appendix F.
Through Lemma 1, we can link $s_{\infty}$ and the reliability gain in Eqs. (48) and (49). In Lemma 1, we mention that $c_{4}\in[0,1]$. The value of $c_{4}$ is determined by the form of $\Psi(\gamma)$. When $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}$ exists and is greater than 0, we have $c_{4}=\frac{1}{2}$. This is because, in this case, we find that $\lim\limits_{\gamma\to+\infty}\frac{\Psi(\gamma)}{\log_{2}\gamma}=\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}$ is also finite. Thus, we have
$\lim_{\gamma\to+\infty}\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}$ (52)
$=$ $\lim_{\gamma\to+\infty}\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}\lim_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}$
$=$ $\lim_{\gamma\to+\infty}\frac{\Psi(\gamma)}{\log_{2}\gamma}\lim_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}=1.$
For $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}=0$ or $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}=+\infty$, $\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{\Psi(\gamma)}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{\Psi(\gamma)}}$ can be 0, +$\infty$+ ∞, or lies in $(0,+\infty)$.
III-C Gain-Conservation Equation for Reliability, Real-time, and Service-Rate Gains
In this subsection, we will propose the revised gain-conservation equation for low-latency communication systems. In contrast to [30], the revised term in the gain-conservation equation accounts for the influence of random arrivals. Before presenting the revised gain-conservation equation, we first provide the definition of service-rate and real-time gains.
The service-rate gain is defined as $\zeta=\lim\limits_{\gamma\to+\infty}\frac{\partial{\Lambda(\gamma)}}{\partial{\log_{2}\gamma}}$, which is similar to the definition of the service-rate gain in [30]. Note that EC indicates the maximum constant source rate that can be supported by the given service process to satisfy the statistical queueing requirement [13]. Thus, $\zeta$ indicates the rate of increase of throughput with the increase of $\log_{2}\gamma$ in the high-SNR regime. According to L’Hospital’s rule [44], we have
$s_{\infty}=\lim\limits_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}=\lim\limits_{\gamma\to+\infty}\frac{\partial{\Lambda(\gamma)}}{\partial{\log_{2}\gamma}}=\zeta.$ (53)
Therefore, we can use the conclusions from Theorem 5 to characterize the service-rate gain $\zeta$.
For the real-time gain, we adopt the same definition as proposed in [30], which is given by $\tau=-\frac{1}{D_{\rm max}}\frac{\partial\log_{2}\delta}{\partial\log_{2}\gamma}$. $\tau$ represents the decay rate of DVP with the increase of $\log_{2}\gamma$ in the high-SNR regime.
Based on Theorem 5, Theorem 6, and Lemma 1, we can then obtain the gain-conversation equation for low-latency communication systems with random arrivals in the FBL regime. The result is summarized in Theorem 7.
Theorem 7. In the FBL regime under the same assumption used in Theorem 6
  • for a SISO system in an AWGN channel, we have
    $\zeta+c_{4}\varpi=1,\,\text{and}$ (54)
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (55)
    By combing Eqs. (54) and (55), we have
    $(1-\theta N\log_{2}e)\zeta+c_{4}\varpi+\tau=1;$ (56)
  • for a SISO system in a Nakagami-$m$ fading channel with a finite $\varpi$, we have
    $\zeta+c_{4}\varpi=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\},\,\text{and}$ (57)
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (58)
    By combing Eqs. (57) and (58), we have
    $(1-\theta N\log_{2}e)\zeta+c_{4}\varpi+\tau=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\}.$ (59)
    In Eqs. (54), (56), (57), and (59), $c_{4}$ are defined in Eq. (51).
Proof:
According to Eqs. (20) and (21), we obtain
$\ln\delta=-\theta N\Lambda(\gamma)D_{\rm max}.$ (60)
Thus, by taking the derivative on both sides of Eq. (60), we have
$-\frac{1}{D_{\rm max}}\lim_{\gamma\to+\infty}\frac{\partial\log_{2}\delta}{\partial\log_{2}\gamma}=\theta N\log_{2}e\lim_{\gamma\to+\infty}\frac{\partial\Lambda(\gamma)}{\partial\log_{2}\gamma}=\left(\theta N\log_{2}e\right)\zeta.$ (61)
By substituting Eqs. (50) and (61) into Eqs. (48), and (49), we obtain the results shown in Theorem 7. ∎
In Theorem 7, we obtain the revised gain-conservation equation for low-latency communication systems. There are some differences between Eq. (59) and Eq. (4) in [30]. First, the parameter $c_{4}$ is 1 in [30]. This is mainly because [30] lacks a discussion on the condition under which normal approximation can be used for high-SNR asymptotic analysis. Second, there is $\theta N$ in Eq. (59), which does not exist in [30]. This is because the arrival in [30] is deterministic. The randomness of the arrival causes a deterioration of the performance of queueing systems with respect to DVP and QoS-constrained throughput.
Moreover, we can obtain the real-time gain based on Theorem 5 and Eq. (58) for Nakagami-$m$ fading channels as
$\tau=\begin{cases}\theta N\log_{2}e,\quad&1\leq\frac{m}{\theta NT\log_{2}e},\\ \frac{m}{T},\quad&1>\frac{m}{\theta NT\log_{2}e}.\end{cases}$ (62)
Considering the case in which $T=1$, we can treat $\tau$ as a revised version of diversity gain defined in [42]. The diversity gain is defined as the ratio of the outage probability to $\log_{2}\gamma$ with $\gamma\to+\infty$, which is equal to $m$ for the Nakagami-$m$ fading channel. According to L’Hospital’s rule, $\tau$ can be treated as the ratio of DVP to $\log_{2}\gamma$ with $\gamma\to+\infty$. From Eq. (62), if $1\leq\frac{m}{\theta NT\log_{2}e}$, $\theta N\log_{2}e\leq m$. Thus, $\tau\leq m$. The diversity gain proposed in [42] is an upper bound of $\tau$. This is because
$\log_{2}\Pr\{D[n]>D_{\rm max}\}$ $\geq\log_{2}\left(\Pr\left\{R^{\epsilon}_{N}[n]<\frac{a[n]}{D_{\rm max}}\right\}\right)^{D_{\rm max}}$ (63)
$\geq D_{\rm max}\log_{2}\left(\Pr\left\{C[n]<\frac{a[n]}{ND_{\rm max}}\right\}\right).$
Thus, we obtain that $-\frac{1}{D_{\rm max}}\frac{\log_{2}\delta}{\log_{2}\gamma}$ is not larger than the diversity gain proposed in [42].
Based on Theorem 7, we provide the conclusions for the Rayleigh fading channel with and without frequency or spatial diversity without proof in Corollary 2.
Corollary 2. In the FBL regime under the assumption used in Theorem 6
  • for a SISO system in a Rayleigh fading channel, we have
    $\zeta+c_{4}\varpi=\min\left\{1,\frac{1}{\theta NT\log_{2}e}\right\},\,\text{and}$ (64)
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (65)
    By combing Eqs. (64) and (65), we have
    $(1-\theta N\log_{2}e)\zeta+c_{4}\varpi+\tau=\min\left\{1,\frac{1}{\theta NT\log_{2}e}\right\}.$ (66)
  • for a system in Rayleigh fading channels with frequency or spatial diversity $\kappa$, we have
    $\zeta+c_{4}\varpi=\min\left\{1,\frac{\kappa}{\theta NT\log_{2}e}\right\},\,\text{and}$ (67)
    $\tau=\left(\theta N\log_{2}e\right)\zeta.$ (68)
    By combing Eqs. (67) and (68), we have
    $(1-\theta N\log_{2}e)\zeta+c_{4}\varpi+\tau=\min\left\{1,\frac{\kappa}{\theta NT\log_{2}e}\right\}.$ (69)
    In Eqs. (64), (66), (67), and (69), $c_{4}$ is the same as in Eq. (51).
III-D $\epsilon-\chi$ Tradeoff in the AWGN Channel
In this subsection, we will take AWGN as an example to characterize the $\epsilon-\chi$ tradeoff through an inequality. The AWGN channel model is widely used in wired communications and satellite communications. With the rapid development of deep learning technology, the communication between processing units, which usually belongs to wired communication, has also raised much attention from both academia and industry. Thus, characterizing the performance limit of the AWGN channel in the FBL regime is important. Different from the discussion in Subsections III-A, III-B, and III-C, we do not use the assumptions shown in Theorem 1 in this subsection.
For the AWGN channel, the DVP is equivalent to QVP since the service rate is fixed given $N$, $\epsilon$, and $\gamma$. Thus, we can choose to use QVP to represent the delay performance of the communication system in the AWGN channel. Similarly, we use $\epsilon$ to evaluate the reliability of the system. Based on the above considerations, we obtain the $\epsilon-\chi$ tradeoff in an inequality in Corollary 3. With some abuse of notations, we let $C=\log_{2}(1+\gamma)$ and $V=\log_{2}^{2}e\left(1-\frac{1}{(1+\gamma)^{2}}\right)$ in Corollary 3.
Corollary 3. Given $\gamma$, $N$, a large $L$, and the requirement $\chi\leq\chi_{\rm th}$, we have
${\color[rgb]{0,0,0}\epsilon\geq Q\left(\frac{NC+\frac{\log_{2}N}{2}+G(N,\gamma,\epsilon)-\alpha_{A}\left(-\frac{1}{L}\ln\chi_{\rm th}\right)}{\sqrt{NV}}\right).}$ (70)
Proof:
See Appendix G.
Corollary 3 presents the relationships between $\epsilon$ and $\chi$. In practical applications, we can use these relationships to determine the proper choice of $\epsilon$ with the given $\chi_{\rm th}$. For the fading channel, we can conduct a similar analysis. However, it is not trivial because it is hard to obtain the analytical expression of $\theta$. With the help of approximation methods and under some assumptions, one may obtain the analytical or even closed-form approximation of $\theta$ [45]. Subsequently, a similar expression to Eq. (70) can be derived, which is beyond the scope of this paper.
IV Effective Capacity Approximation with Normal Approximation
In this section, we present an EC approximation method for low-latency communication systems. The closed-form expression of the approximation characterizes the influence of the channel distribution, error probability, and blocklength on the EC. As shown in Eqs. (19) and (21), there are limitations on the use of LDT to determine the QVP and DVP in URLLC scenarios. However, EC is widely used in the works for URLLC scenarios [26, 27, 24, 25] to evaluate the QoS-constrained throughput. In these studies, the value of the QoS exponent $\theta$ is given to represent the QoS requirement, while the EC is adopted to characterize the throughput under the QoS requirement. Some scheduling schemes in URLLC were proposed to maximize the EC in these works. Therefore, deriving the closed-form expression of EC is crucial for further optimization. However, it is challenging to derive a closed-form expression of the EC, complicating throughput analysis and the comparison between different scheduling schemes. To address this issue, we propose a Laplace’s-method-based approximation for EC in this section. In this subsection, we set $T\geq 1$, which covers the general low-latency scenarios. Note that we use normal approximation to approximate the maximum coding rate in this section. Thus, the proposed approximation can serve as an efficient EC approximation method in scenarios where normal approximation achieves satisfactory accuracy.
IV-A EC Approximation for Fixed-Power Allocation
In this subsection, we start with the simple case in which the transmitted power is constant in each time slot. For low-latency scenarios, the threshold $L$ and $D_{\rm max}$ are usually small due to the stringent QoS requirement [22]. Thus, to achieve an extremely small QVP or DVP, the value of the QoS exponent $\theta$ or the revised QoS exponent $\theta N$ needs to be large. By this means, $\theta NT$ should have a large value. Based on this thought, we start our analysis of EC with the fixed-power allocation in Lemma 2. Note that $\ell=\left(\frac{Q^{-1}(\epsilon)}{\sqrt{N}}\right)^{2}$ as defined before Theorem 4.
Lemma 2. For a low-latency communication system adopting fixed-power allocation, EC is approximated as
$\alpha_{S}(\theta)=Nr(x^{*})-\frac{\ln f(x^{*})+\frac{1}{2}\ln 2\pi-\frac{1}{2}\ln|r^{\prime\prime}(x^{*})|-\frac{1}{2}\ln\theta NT}{\theta T}+O\left(\frac{1}{\theta^{2}T^{2}N}\right),$ (71)
where $r(x)$ is defined in Eq. (40) and $x^{*}$ is given by
$x^{*}=\frac{1}{\gamma}\left(\sqrt{\frac{1+\sqrt{1+4\ell}}{2}}-1\right).$ (72)
Proof:
See Appendix H.
In Lemma 2, we provide an approximation of EC for the fixed-power allocation, which provides a closed-form expression of EC with arbitrary fading channels. Besides, the approximation error is $O\left(\frac{1}{\theta^{2}T^{2}N}\right)$. Thus, as we mentioned before Lemma 2, this approximation performs better in a low-latency communication system, which we will validate by numerical calculation in Section V.
IV-B EC Approximation with Channel-Gain-Based Power Allocation
In this subsection, we will generalize the proposed EC approximation method for the fixed-power allocation scheme to certain types of channel-gain-based power allocation schemes. The conditions for using the proposed approximation are presented. Moreover, we discuss the applicability of the proposed approximation for some classical channel-gain-based power allocation schemes.
Based on Lemma 2, we can extend the approximation to more general power allocation schemes. We summarize this generalization in Theorem 8.
Theorem 8. For a low-latency communication system which adopts a channel-gain-based power allocation scheme $\Xi(x)$ satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ with channel distribution $f(\cdot)$, EC is approximated as
$\alpha_{S}(\theta)=Nr(x^{*})-\frac{\ln f(x^{*})+\frac{1}{2}\ln 2\pi-\frac{1}{2}\ln|r^{\prime\prime}(x^{*})|-\frac{1}{2}\ln\theta NT}{\theta T}+O\left(\frac{1}{\theta^{2}T^{2}N}\right),$ (73)
where $r(x)$ is defined in Eq. (40) and $x^{*}$ is given by
$\Xi(x^{*})x^{*}=\sqrt{\frac{1+\sqrt{1+4\ell}}{2}}-1.$ (74)
In Eq. (74), $\ell=\left(\frac{Q^{-1}(\epsilon)}{\sqrt{N}}\right)^{2}$.
Proof:
See Appendix I.
In Theorem 8, we propose the EC approximation for a specific kind of power allocation scheme. For the policies satisfying $\frac{{\rm d}\Xi(x)}{{\rm d}x}\geq 0$, $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ holds. $\frac{{\rm d}\Xi(x)}{{\rm d}x}\geq 0$ means that the allocated power does not decrease with the increasing channel power gain. This condition can be met with some power allocation schemes, one of which is the well-known water-filling scheme [46]. Furthermore, when $\frac{{\rm d}\Xi(x)}{{\rm d}x}\geq 0$, $\Xi(x)x$ is a monotonic function of $x$. Thus, the solution of Eq. (42) can be obtained directly via binary search.
A power-control policy that maximizes the effective capacity under the average power constraint is proposed in [47]. This policy also satisfies the constraint presented in Theorem 8. Let $\Xi_{1}(x)$ denote the power-control policy proposed in [47], which can be expressed as
$\Xi_{1}(x)=\begin{cases}\frac{1}{a_{1}^{\frac{1}{a_{2}+1}}(\gamma x)^{\frac{a_{2}}{a_{2}+1}}}-\frac{1}{\gamma x},\quad&x\geq\frac{a_{1}}{\gamma},\\ 0,\quad&x<\frac{a_{1}}{\gamma},\end{cases}$ (75)
where $a_{1}$ and $a_{2}$ are positive constants. $a_{1}$ is related to the average power constraint, while $a_{2}$ is a linear and increasing function of $\theta$. For $x\geq\frac{a_{1}}{\gamma}$, we have
$\Xi_{1}(x)+\frac{{\rm d}\Xi_{1}(x)}{{\rm d}x}x$ $=\frac{1}{a_{1}^{\frac{1}{a_{2}+1}}(\gamma x)^{\frac{a_{2}}{a_{2}+1}}}-\frac{1}{\gamma x}+x\left(-\frac{a_{2}}{a_{2}+1}\frac{1}{a_{1}^{\frac{1}{a_{2}+1}}\gamma^{\frac{a_{2}}{a_{2}+1}}}x^{-\frac{a_{2}}{a_{2}+1}-1}+\frac{1}{\gamma}\frac{1}{x^{2}}\right)$ (76)
$=\frac{1-\frac{a_{2}}{a_{2}+1}}{a_{1}^{\frac{1}{a_{2}+1}}\gamma^{\frac{a_{2}}{a_{2}+1}}}x^{-\frac{a_{2}}{a_{2}+1}}.$
Since $\frac{a_{2}}{a_{2}+1}<1$, it follows that $\Xi_{1}(x)+\frac{{\rm d}\Xi_{1}(x)}{{\rm d}x}x\geq 0$ for $x>0$, which means the power-control policy proposed in [47] satisfies the condition of Theorem 8. There are two special cases to consider. When $a_{2}$ approaches 0, the power-control policy proposed in [47] degenerates to the water-filling policy, which has been discussed above. When $a_{2}$ approaches infinity, the power-control policy proposed in [47] degenerates to channel inversion policy, with which the rate is constant over all channel states. In this case, EC is trivial, as the rate remains fixed.
Although the approximation provided in Theorem 8 can be applied to the policies satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$, the accuracy of this approximation can be further enhanced through additional optimization tailored to the specific properties of each policy. For example, in the water-filling scheme, there is a channel-gain threshold $h_{\rm th}$. If the channel gain falls below this threshold, the allocated power is 0, which means the transmission rate is constant in the FBL regime when $h[n]<h_{\rm th}$. [^9] The true rate should be zero when $\gamma=0$. However, since we use the normal approximation as we mentioned at the beginning of this section, we let the rate be equal to $\frac{\log_{2}N}{2}$ when $\gamma=0$ according to (9). This setting can be easily adjusted by setting the rate to 0 when $\gamma=0$, which does not affect the accuracy of this approximation. Let $\xi(\theta)=e^{-\theta T\frac{\log_{2}(N)}{2}}\int_{0}^{h_{\rm th}}f(x){\rm d}x$. Based on the above analysis and Theorem 8, EC can be approximated as
$\alpha_{S}(\theta)=-\frac{1}{\theta N}\ln\left(\xi(\theta)+\sqrt{\frac{2\pi}{\theta NT|r^{\prime\prime}(x^{*})|}}f(x^{*})e^{-\theta NTr(x^{*})}\left(1+O\left((\theta NT)^{-1}\right)\right)\right).$ (77)
For those policies exhibiting truncated characteristics, Eq. (77) provides a more accurate approximation of their EC.
For the low-latency system with the retransmission mechanism, the EC is discussed in [28], which differs from the above analysis without the retransmission mechanism. In these works, a simple ARQ mechanism is assumed to be adopted. With this mechanism, the receiver sends a negative acknowledgment requesting the retransmission of the message in case of an erroneous reception. In Corollary 4, we generalize the EC approximation to the low-latency communication system incorporating this ARQ mechanism.
Corollary 4. For a low-latency system with the above ARQ mechanism and a channel-gain-based power allocation scheme $\Xi(x)$ satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ with channel distribution $f(\cdot)$, EC is approximated as
$\alpha_{S}(\theta)=-\frac{1}{\theta T}\ln\left(\epsilon+(1-\epsilon)\sqrt{\frac{2\pi}{\theta NT|r^{\prime\prime}(x^{*})|}}f(x^{*})e^{-\theta NTr(x^{*})}\left(1+O\left((\theta NT)^{-1}\right)\right)\right),$ (78)
where $r(x)$ is defined in Eq. (40) and $x^{*}$ is given by Eq. (42).
Proof:
See Appendix J.
Figure 4: $s_{\infty}$ in the Rayleigh fading channel.
Figure 5: $s_{\infty}$ in Rayleigh fading channels with $\kappa=2$.
(a) $s_{\infty}$ in the Nakagami-$m$ channel with $m=0.5$.
(b) $s_{\infty}$ in the Nakagami-$m$ channel with $m=2$.
Figure 6: $s_{\infty}$ in the Nakagami-$m$ fading channel.
V Numerical Results
In this section, we conduct numerical calculations to validate the theoretical results and conclusions in this paper. Moreover, some characteristics of the statistical-delay performance of low-latency systems in the FBL regime will be explained more visually by the numerical results.
First, we validate the conclusions of high-SNR slopes of EC in the FBL regime in Figs. 5, 5, and 6. In these figures, we set $N=512$, $\epsilon=10^{-5}$, and $T=1$. In Fig. 5, $\mathbb{E}_{|h|^{2}}\{|h[n]|^{2}\}=1$. In Fig. 5, $f(x)$ satisfies Erlang distribution with $\kappa=2$ and $\mathbb{E}_{|h|^{2}}\{|h[n]|^{2}\}=2$. In Fig. 6, $\Omega=1$. For the Rayleigh fading case, $\frac{1}{T\log_{2}e}\approx 0.6931$. From Fig. 5, we find that the slopes of $\Lambda(\gamma)$ for $\theta N\leq 0.6931$, i.e. $\theta\leq 0.0014$, are nearly identical and close to 1, while the slopes of $\Lambda(\gamma)$ for $\theta>0.0014$ are smaller than 1. Besides, for those $\theta>0.0014$, a larger $\theta$ results in a smaller slope of $\Lambda(\gamma)$. The slope of $\Lambda(\gamma)$ for $\theta=0.004$ is $0.338$ in Fig. 5, while $\frac{1}{\theta NT\log_{2}e}$ with $\theta=0.004$ is 0.3385. The above results validate the accuracy of Corollary 1. Then we analyze the numerical results of the case in which $f(x)$ satisfies the Erlang distribution. In Fig. 5, $\frac{\kappa}{T\log_{2}e}\approx 1.3863$. We can find that with $\theta=0.002$, $\theta N=1.024$ is larger than $\frac{1}{T\log_{2}e}\approx 0.6931$ in Fig. 5 while it is smaller than $\frac{\kappa}{T\log_{2}e}\approx 1.3863$ in Fig. 5. Accordingly, the slope of $\Lambda(\gamma)$ for $\theta=0.002$ in Fig. 5 is different from the slope of ergodic capacity, while the slope of $\Lambda(\gamma)$ for $\theta=0.002$ is nearly identical to the slope of ergodic capacity in Fig. 5. This result reveals the difference between the performance of $s_{\infty}$ with and without the spatial or frequency diversity in Rayleigh fading channels, which also validates Corollary 1.
Figure 7: EC approximation for the fixed-power allocation scheme.
In Fig. 6, we present the numerical results of $s_{\infty}$ in the Nakagami-$m$ fading channel. In Fig. 6(a), $\frac{m}{TN\log_{2}e}\approx 0.3466$, while in Fig. 6(b), $\frac{m}{TN\log_{2}e}\approx 1.3836$. The slopes of $\Lambda(\gamma)$ in Figs. 6(a) and 6(b) validates the accuracy of Theorem 5. Besides, by comparing the results in Figs. 5 and 6(b), we find that the slopes of $\Lambda(\gamma)$ is nearly identical with a same $\theta$. For example, when $\theta=0.004$, i.e., $\theta N=2.0480$, the slope of $\Lambda(\gamma)$ in Fig. 5 is $0.6731$, while the slope of $\Lambda(\gamma)$ in Fig. 6(b) is $0.6704$. This result also validates the accuracy of Theorem 5 and Corollary 1, since $s_{\infty}$ with $\kappa=2$ and $m=2$ are identical as shown in Theorem 5 and Corollary 1.
Figure 8: EC approximation for the water-filling policy.
Furthermore, we validate the accuracy of the approximation for EC by numerical calculations. First, we present the results of the EC approximation with the fixed-power allocation in Fig. 7. In this figure, $\gamma=10$ and $\epsilon=10^{-3}$. The channel model is set as the Rayleigh fading channel with $\mathbb{E}_{|h|^{2}}\{|h[n]|^{2}\}=1$. The numerical result in this figure is calculated based on the definition of the EC in Eq. (16), while the approximation result is obtained from Eq. (71). From Fig. 7, we find that the approximation results match well with the numerical results. The difference between the numerical and approximation results with $N=100$ and $T=1$ is slightly more evident when $\theta$ is small compared to the other three pairs. This is reasonable since the approximation error of Laplace’s method is $O\left(\frac{1}{\theta^{2}T^{2}N}\right)$. From these results, we see that the approximation in Lemma 2 satisfactorily approximates the EC in the low-latency scenarios.
Fig. 8 presents the result of EC approximation with the water-filling (WF) policy [46]. To test the performance of the EC approximation in different channel models, in Fig. 8 the channel gain satisfies $|h[n]|^{2}=|h_{1}[n]|^{2}+|h_{2}[n]|^{2}$, where $|h_{i}[n]|^{2}$ follows an exponential distribution with expectation 1. The approximation for the WF policy we adopted here is based on Eq. (77). From Fig. 8, we can see that the approximation and numerical results of the WF policy match well, which validates our analysis in Section III-C. Note that, for the water-filling policy with the average SNR $\mathbb{E}_{|h|^{2}}\{\gamma\}=10$, [^10] With some abuse of notations, we use $\gamma$ to denote the SNR of the water-filling policy, which should be a random variable related to $|h[n]|^{2}$. $\mathbb{E}_{{|h|^{2}}}\{s[n]\}\approx 3932$, while for the fixed-power allocation policy with $\gamma=10$, $\mathbb{E}_{|h|^{2}}\{s[n]\}\approx 3923$. This result indicates that the transmission optimization based on the infinite blocklength assumption can provide help for improving the average transmission rate in the FBL regime under some parameter settings. However, the result in Fig. 8 shows that, though the WF policy has a larger average transmission rate compared with the fixed-power allocation scheme, it does not improve the EC. Since the EC represents the maximum throughput under the QVP or DVP constraints, the poorer performance of the WF policy on the EC demonstrates that there is a difference between maximizing the average transmission rate and maximizing the throughput under statistical QoS constraints.
VI Conclusion
In this paper, we mainly focused on characterizing the reliability-latency-rate tradeoff for low-latency systems with FBC. Based on the previous conclusions on the gain-conservation property for low-latency systems with deterministic arrival, we defined the reliability gain, real-time gain, and service-rate gain for the low-latency systems with random arrivals. EB and EC were adopted to establish the connection between the gain-conservation equation and the influence of random arrivals. By considering transmissions over an AWGN channel and a Nakagami-$m$ fading channel, we derived the service-rate gains in wired and wireless communication systems. Then, the revised gain-conservation equations were derived by taking service-rate, reliability, and real-time gains into account. Moreover, an EC approximation method was proposed to obtain the analytical expression of EC for some channel-gain-based scheduling policies in the FBL regime. The conclusions proposed in this paper characterize the fundamental tradeoff between the reliability, latency, and throughput for low-latency systems with FBC, which will provide instructions on the design of efficient scheduling schemes for the low-latency systems. Future promising research directions include obtaining more accurate approximations of QVP and DVP with the small threshold, relaxing the assumptions of the proposed analysis through tighter bounds or exact expression of the FBL coding rate, and designing efficient scheduling schemes based on the revised gain-conservation equations and proposed EC approximation method for low-latency communication systems.
Appendix A Proof of Theorem 2
According to Eqs. (30) and (34), we have
${\color[rgb]{0,0,0}g_{u}(|h|^{2}\gamma,\epsilon)=\frac{c_{1}}{V\left(|h|^{2}\gamma\right)}+\frac{3}{2}\log_{2}V\left(|h|^{2}\gamma\right)-\log_{2}K\left(|h|^{2}\gamma\right),}\vspace{-1mm}$ (79)
and
$g_{l}(N,|h|^{2}\gamma,\epsilon)$ $=-\frac{1}{2}\log_{2}N+\log_{2}\frac{K\left(|h|^{2}\gamma\right)}{c_{2}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}-\log_{2}\left[2\left(\frac{\ln 2}{\sqrt{2\pi}}+\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)\right]$ (80)
$\quad\quad\quad+\sqrt{NV\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{N}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon)\right],$
Since $|h|^{2}$ and $\gamma$ are independent, the product $|h|^{2}\gamma$ can be zero, finite, or approach positive infinity. We need to consider all possible cases of $|h|^{2}\gamma$. The essence of this proof is to demonstrate that $G(N,|h|^{2}\gamma,\epsilon)$ is bounded with the arbitrary value of $|h|^{2}\gamma$. We divide the proof into four parts. We let $\vartheta>0$ denote an arbitrarily small constant. The first part is the proof with $|h|^{2}\gamma\to 0$, i.e., $|h|^{2}\gamma\in[0,\vartheta)$. The second part is the proof with a bounded $\Psi(\gamma)$ and $|h|^{2}\gamma\in[\vartheta,+\infty)$. The third part is the proof with $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$ and $|h|^{2}\gamma\in[\vartheta,+\infty)$. Finally, we will summarize the proof in the fourth subsection. To use these bounds, we assume that
$N>\max\left\{\left(\frac{2K(|h|^{2}\gamma)}{(1-\epsilon)V^{\frac{3}{2}}(|h|^{2}\gamma)}\right)^{2},\left(\frac{2K(|h|^{2}\gamma)}{\epsilon V^{\frac{3}{2}}(|h|^{2}\gamma)}\right)^{2}\right\}.$ (81)
The validity of Eq. (81) is provided in the fourth part.
A-A Proof for $|h|^{2}\gamma\to 0$, i.e., $|h|^{2}\gamma\in[0,\vartheta)$
The analysis of $|h|^{2}\gamma\to 0$ is based on the physical significance. The proof in this subsection is to demonstrate that $G(N,|h|^{2}\gamma,\epsilon)$ is bounded when $|h|^{2}\gamma\to 0$. With $|h|^{2}\gamma\in\to 0$, it is obvious that $\tilde{R}^{\epsilon}_{N}$ is bounded according to Eq. (9). Since we can not obtain an unbounded rate with a finite received SNR, the coding rate $\tilde{R}^{\epsilon}_{N}+G(N,|h|^{2}\gamma,\epsilon)$ is also bounded as $|h|^{2}\gamma\to 0$. Thus, $G(N,|h|^{2}\gamma,\epsilon)$ must be bounded when $|h|^{2}\gamma\to 0$. This result implies $|G(N,|h|^{2}\gamma,\epsilon)|\leq u_{1}$ for $|h|^{2}\gamma\to 0$, where $u_{1}$ is a positive constant.
A-B Proof for a Bounded $\Psi(\gamma)$ and $|h|^{2}\gamma\in[\vartheta,+\infty)$
We use the upper and lower bounds to prove that $G(N,|h|^{2}\gamma,\epsilon)$ is bounded as $|h|^{2}\gamma\to+\infty$. We start the analysis with the upper bound $g_{u}(|h|^{2}\gamma,\epsilon)$. According to Eq. (31), $K\left(|h|^{2}\gamma\right)$ is finite for a given finite $|h|^{2}\gamma\in[\vartheta,+\infty)$. As $|h|^{2}\gamma\to+\infty$, $\frac{\gamma}{1+\gamma}\to 1$ and
$\lim_{|h|^{2}\gamma\to+\infty}\mathbb{E}_{z}\left\{\left|z^{2}-\frac{2}{\sqrt{|h|^{2}\gamma}}z-1\right|^{3}\right\}<+\infty.$ (82)
Thus, $K\left(|h|^{2}\gamma\right)$ is also bounded when $|h|^{2}\gamma\in[\vartheta,+\infty)$.
Moreover, we have $\left(1-\frac{1}{(1+\vartheta)^{2}}\right)\left(\log_{2}e\right)^{2}\leq V\left(|h|^{2}\gamma\right)\leq\left(\log_{2}e\right)^{2}$ when $|h|^{2}\gamma\in[\vartheta,+\infty)$. Thus, we have obtained that $K\left(|h|^{2}\gamma\right)$ and $V\left(|h|^{2}\gamma\right)$ is bounded for $|h|^{2}\gamma\in[\vartheta,+\infty)$.
Then, we will analyze $b_{1}$. First, we want to show that $b_{1}<1-\epsilon$, i.e.,
$\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)\sqrt{\left(\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)(1-\epsilon)}\right)^{2}+1}}<1-\epsilon.$ (83)
By manipulating Eq. (83), it is equivalent to show that the following inequality holds.
${\color[rgb]{0,0,0}\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{(2K\left(|h|^{2}\gamma\right))^{2}+V^{3}\left(|h|^{2}\gamma\right)(1-\epsilon)^{2}}}<1.}$ (84)
Eq. (84) is true since $\left(1-\frac{1}{(1+\vartheta)^{2}}\right)\left(\log_{2}e\right)^{2}\leq V\left(|h|^{2}\gamma\right)$ with $|h|^{2}\gamma\in[\vartheta,+\infty)$. Thus, $01-\epsilon-b_{1}>$ 0, which indicates $\min_{y\in\left[1-\epsilon-b_{1},1-\epsilon\right]}\frac{{\rm d}Q^{-1}(y)}{{\rm d}y}$ is finite. Therefore, $c_{1}$ is bounded for $|h|^{2}\gamma\in[\vartheta,+\infty)$ according to Eq. (32).
We have proved that for $|h|^{2}\gamma\in[\vartheta,+\infty)$, every term of $g_{u}\left(|h|^{2}\gamma,\epsilon\right)$ shown in Eq. (79) is finite. Thus, $|g_{u}\left(|h|^{2}\gamma,\epsilon\right)|\leq u_{2}$ for $|h|^{2}\gamma\in[\vartheta,+\infty)$, where $u_{2}$ is a positive constant.
Then, we will analyze $g_{l}(N,|h|^{2}\gamma,\epsilon)$. We consider two cases. Similar to the analysis of $g_{u}\left(|h|^{2}\gamma,\epsilon\right)$, every term of $g_{l}(N,|h|^{2}\gamma,\epsilon)$ is finite for $|h|^{2}\gamma\in[\vartheta,+\infty)$. This result implies $\left|g_{l}(N,|h|^{2}\gamma,\epsilon)\right|\leq u_{3}$, where $u_{3}$ is a positive constant.
In summary, we have proved that $|g_{u}\left(|h|^{2}\gamma,\epsilon\right)|$ and $|g_{l}(N,|h|^{2}\gamma,\epsilon)|$ are all bounded for $|h|^{2}\gamma\in[\vartheta,+\infty)$ and a bounded $\Psi(\gamma)$. Since $g_{l}(N,|h|^{2}\gamma,\epsilon)\leq G(N,|h|^{2}\gamma,\epsilon)\leq g_{u}(|h|^{2}\gamma,\epsilon)$, we have $|G(N,|h|^{2}\gamma,\epsilon)|\leq\max\{u_{2},u_{3}\}$ for $|h|^{2}\gamma\in[\vartheta,+\infty)$.
Combining the results from the last subsection with this subsection, we have proved that
$|G(N,|h|^{2}\gamma,\epsilon)|\leq\max\{u_{1},u_{2},u_{3}\},\quad|h|^{2}\gamma\in[0,+\infty),$ (85)
holds with a given $\epsilon$, a bounded $\Psi(\gamma)$, and a given $N$ satisfying Eq. (81). With a bounded $\Psi(\gamma)$, Eq. (85) ensures Eq. (22) is satisfied. Thus, to satisfy the condition in Eq. (22) with a given $\epsilon$ and a bounded $\Psi(\gamma)$, we only need to ensure that Eq. (81) is met.
A-C Proof for $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$ and $|h|^{2}\gamma\in[\vartheta,+\infty)$
Since the upper bound is irrelevant to $N$, the only difference between this part and the last part lies in the analysis of the lower bound $g_{l}(N,|h|^{2}\gamma,\epsilon)$. For $|h|^{2}\gamma\in[0,\vartheta)$, the proof is the same as the proof in Subsection A. Thus, we only need to focus on $|h|^{2}\gamma\in[\vartheta,+\infty)$. With $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, there are two unbounded terms in $g_{l}(N,|h|^{2}\gamma,\epsilon)$. The unbounded terms are $-\frac{1}{2}\log_{2}\Psi(\gamma)$ and $\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon(\gamma))\right]$. Since $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, we have
$\lim_{\gamma\to+\infty}\frac{-\frac{1}{2}\log_{2}\Psi(\gamma)}{\Psi(\gamma)}=0.$ (86)
Then, we turn to the term $\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon)\right]$. By performing a Taylor-series expansion of the inverse Gaussian Q-function around $1-\epsilon$, we obtain
$\lim_{\gamma\to+\infty}\frac{\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon)\right]}{\Psi(\gamma)}$ (87)
$=$ $\lim_{\gamma\to+\infty}\frac{\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon\right)+Q^{-1}(\epsilon)+O\left(\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)\right]}{\Psi(\gamma)}$
$=$ $\lim_{\gamma\to+\infty}\frac{\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}O\left(\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)}{\Psi(\gamma)}$
$=$ $0$
Therefore, for the lower bound $g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))$, any $\Psi(\gamma)$ that satisfies $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$ can ensure $\lim\limits_{\gamma\to+\infty}\frac{g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}=0$. Since the upper bound is bounded, we have $\lim\limits_{\gamma\to+\infty}\frac{g_{u}(|h|^{2}\gamma,\epsilon)}{\Psi(\gamma)}=0$. Since we use the upper and lower bounds, we need to ensure that Eq. (81) holds. Thus, we have proved that, with a $\Psi(\gamma)$ satisfying Eq. (81) and $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, Eq. (22) holds.
A-D Summary of the Proof
From the above proof, we find that whether $\Psi(\gamma)$ is bounded or $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, Eq. (81) is enough to ensure that Eq. (22) holds with a given $\epsilon$. We let
$\varsigma_{1}=\max_{x\geq 0}\left\{\left(\frac{2K(x)}{(1-\epsilon)V^{\frac{3}{2}}(x)}\right)^{2},\left(\frac{2K(x)}{\epsilon V^{\frac{3}{2}}(x)}\right)^{2}\right\}.$ (88)
Therefore, to ensure that Eq. (81) holds with a given $\epsilon$, we can choose a $\Psi(\gamma)$ satisfying
${\color[rgb]{0,0,0}\Psi(\gamma)>\varsigma_{1}.}$ (89)
To ensure the validity of Eq. (89), we have to prove that $\varsigma_{1}$ is bounded. It is equivalent to proving that $\left(\frac{2K(x)}{V^{\frac{3}{2}}(x)}\right)^{2}$ is bounded on $[0,+\infty)$. For $\left(\frac{2K(x)}{V^{\frac{3}{2}}(x)}\right)^{2}$, its value is bounded for $x\to+\infty$ and finite $x$ as we analyzed above. For $x\to 0$, we have
$\lim_{x\to 0}\frac{2K(x)}{V^{\frac{3}{2}}(x)}=$ $\frac{2c_{0}}{(\log_{2}e)^{3}}\lim_{x\to 0}\frac{x^{3}\mathbb{E}_{z}\left\{\left|z^{2}-\frac{2}{\sqrt{x}}z-1\right|^{3}\right\}}{(1+x)^{3}\left(1-\frac{1}{(1+x)^{2}}\right)^{\frac{3}{2}}}$ (90)
$=$ $\frac{2c_{0}}{(\log_{2}e)^{3}}\lim_{x\to 0}\frac{x^{\frac{3}{2}}}{(x+2)^{\frac{3}{2}}}\left(u_{4}x^{-\frac{3}{2}}+o\left(x^{-\frac{3}{2}}\right)\right)=u_{5},$
where $u_{4}$ and $u_{5}$ are finite constant. $o\left(x^{-\frac{3}{2}}\right)$ satisfies $\lim\limits_{x\to 0}\frac{o\left(x^{-\frac{3}{2}}\right)}{x^{-\frac{3}{2}}}=0$. Thus, we prove that $\varsigma_{1}$ is bounded. The proof is completed.
Appendix B Proof of Theorem 3
For finite $|h|^{2}\gamma$, the proof is the same as that in Appendix A, in which $g_{u}(|h|^{2}\gamma,\epsilon(\gamma))$ and $g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))$ are all bounded. Thus, in the following proof, we mainly focus on the analysis for $|h|^{2}\gamma\to+\infty$.
Note that we use the upper and lower bounds in the following proof. Thus, this result is under the assumption that
$\Psi(\gamma)>\max\left\{\left(\frac{2K(|h|^{2}\gamma)}{(1-\epsilon(\gamma))V^{\frac{3}{2}}(|h|^{2}\gamma)}\right)^{2},\left(\frac{2K(|h|^{2}\gamma)}{\epsilon(\gamma)V^{\frac{3}{2}}(|h|^{2}\gamma)}\right)^{2}\right\}.$ (91)
Since $\epsilon(\gamma)\in(0,0.5]$, we have $1-\epsilon(\gamma)\geq\epsilon(\gamma)$. Thus, Eq. (91) is equivalent to
${\color[rgb]{0,0,0}\Psi(\gamma)>\left(\frac{2K(|h|^{2}\gamma)}{\epsilon(\gamma)V^{\frac{3}{2}}(|h|^{2}\gamma)}\right)^{2}.}$ (92)
Let $\varsigma_{2}=4\max_{x\geq 0}\left\{\frac{K^{2}(x)}{V^{3}(x)}\right\}$. Similar to Eq. (90), $\varsigma_{2}$ is bounded. Thus, we can let $\Psi(\gamma)\epsilon^{2}(\gamma)>\varsigma_{2}$ to ensure Eq. (91) is satisfied. We let $\Psi(\gamma)\epsilon^{2}(\gamma)=(1+\omega)\varsigma_{2}$, where $\omega>0$ is an arbitrarily positive constant. Then we have
${\color[rgb]{0,0,0}\epsilon(\gamma)=\sqrt{\frac{(1+\omega)\varsigma_{2}}{\Psi(\gamma)}}.}$ (93)
Eq. (93) will be used below.
B-A Analysis of $g_{u}(|h|^{2}\gamma,\epsilon(\gamma))$ for $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$ and $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)\to+\infty$
Given that $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$, we have $1-\epsilon(\gamma)\to 1$ for $\gamma\to+\infty$. Thus, according to Eq. (32), $c_{1}\to+\infty$ for $\gamma\to+\infty$ since $\frac{dQ^{-1}(y)}{dy}\Big{|}_{y=1-\epsilon(\gamma)}\to-\infty$. Except for $c_{1}$, the other terms in $g_{u}(|h|^{2}\gamma,\epsilon(\gamma))$ remains bounded. Thus, to ensure that there exists a finite $\nu$ such that $\lim\limits_{\gamma\to+\infty}\left|\frac{g_{u}(|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}\right|\leq\nu$ holds, it is equivalent to requiring
${\color[rgb]{0,0,0}\lim_{\gamma\to+\infty}\frac{\frac{dQ^{-1}(1-\epsilon(\gamma))}{dy}}{\Psi(\gamma)}<+\infty.}$ (94)
By substituting the derivative of inverse Gaussian Q-function $\frac{{\rm d}Q^{-1}(y)}{{\rm d}y}=\left[-\frac{1}{\sqrt{2\pi}}\exp\left(-\left(\frac{Q^{-1}(y)}{2}\right)^{2}\right)\right]^{-1}$ into Eq. (94), it is equivalent to requiring
${\color[rgb]{0,0,0}\lim_{\gamma\to+\infty}\frac{e^{\left(\frac{Q^{-1}(1-\epsilon(\gamma))}{2}\right)^{2}}}{\Psi(\gamma)}=\lim_{\gamma\to+\infty}\frac{e^{\left(\frac{Q^{-1}(\epsilon(\gamma))}{2}\right)^{2}}}{\Psi(\gamma)}<+\infty.}$ (95)
B-B Analysis of $g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))$ for $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$ and $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)\to+\infty$
Given that $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$ and $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)\to+\infty$, the two terms in Eq. (80) are not bounded, which are the same as those mentioned in Appendix A. For $-\frac{1}{2}\log_{2}\Psi(\gamma)$, it follows Eq. (86) without changes. For simplicity, we let
${\color[rgb]{0,0,0}H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))=\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon(\gamma))\right]}$ (96)
denote another unbounded term. Since $Q^{-1}(x)$ is a decreasing function and $\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}>0$, we have
${\color[rgb]{0,0,0}Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon(\gamma))\leq Q^{-1}\left(1-\epsilon(\gamma)\right)+Q^{-1}(\epsilon(\gamma))=0.}$ (97)
Based on Eq. (97), we have
${\color[rgb]{0,0,0}\lim_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}\leq 0.}$ (98)
We have proved the upper bound of $\lim\limits_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}$ is finite. Then, we demonstrate the condition with which the lower bound of $\lim\limits_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}$ is finite. According to Eq. (93) and the definition of $\varsigma_{2}$, we have
$Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)$ $\geq Q^{-1}\left(1-\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)$ (99)
$=-Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right).$
Then we have
$\lim_{\gamma\to+\infty}\frac{\sqrt{\Psi(\gamma)V\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon(\gamma)+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{\Psi(\gamma)}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon(\gamma))\right]}{\Psi(\gamma)}$ (100)
$\geq$ $\lim_{\gamma\to+\infty}\frac{-Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)+Q^{-1}(\epsilon(\gamma))}{\sqrt{\Psi(\gamma)}}.$
We will give the proof of $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}=0$ in Appendix C. Thus, $\lim\limits_{\gamma\to+\infty}\frac{H(\gamma,|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}<+\infty$ holds if
${\color[rgb]{0,0,0}\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{\Psi(\gamma)}}<+\infty.}$ (101)
Thus, if Eq. (101) holds, we have $\lim\limits_{\gamma\to+\infty}\left|\frac{g_{l}(\Psi(\gamma),|h|^{2}\gamma,\epsilon(\gamma))}{\Psi(\gamma)}\right|\leq\nu$.
Therefore, if Eqs. (95), (101), and $\Psi(\gamma)\epsilon^{2}(\gamma)>\varsigma_{2}$ hold, the condition shown in Eq. (22) is satisfied. The proof is completed.
Appendix C Proof of $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}=0$
Since $\Psi(\gamma)\epsilon^{2}(\gamma)=(1+\omega)\varsigma_{2}$, we have $\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)<\epsilon(\gamma)<\frac{1}{2}$. Thus, we obtain $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}\geq 0$. To adopt the squeeze theorem, we then need to demonstrate that $\lim\limits_{\gamma\to+\infty}\frac{Q^{-1}\left(\sqrt{\frac{\varsigma_{2}}{\Psi(\gamma)}}\left(\sqrt{1+\omega}-1\right)\right)}{\sqrt{\Psi(\gamma)}}\leq 0$. For simplicity, we let $\omega_{1}=\sqrt{\varsigma_{2}}\left(\sqrt{1+\omega}-1\right)$ in the following proof. We let $Q^{-1}\left(v\right)=x$, which can also be written as $v=Q(x)$. Based on Mills’ ratio, for $x>0$, we obtain that
${\color[rgb]{0,0,0}\frac{1}{\sqrt{2\pi}}\frac{x}{1+x^{2}}e^{-\frac{x^{2}}{2}}\leq v\leq\frac{1}{\sqrt{2\pi}}\frac{1}{x}e^{-\frac{x^{2}}{2}}.}$ (102)
Then, we have
${\color[rgb]{0,0,0}x^{2}+\ln x^{2}\leq-2\ln\left(\sqrt{2\pi}v\right).}$ (103)
Let us introduce the Lambert W function $W(\cdot)$. $W(v)$ is the solution of $ye^{y}=v$. Based on $W(\cdot)$, we can transform Eq. (103) into
${\color[rgb]{0,0,0}-\sqrt{W\left(\frac{v^{-2}}{2\pi}\right)}\leq x\leq\sqrt{W\left(\frac{v^{-2}}{2\pi}\right)}.}$ (104)
Since $\frac{\omega_{1}}{\sqrt{\Psi(\gamma)}}<\frac{1}{2}$, we have
${\color[rgb]{0,0,0}0<Q^{-1}\left(\frac{\omega_{1}}{\sqrt{\Psi(\gamma)}}\right)<\sqrt{W\left(\frac{\Psi(\gamma)}{2\pi\omega_{1}^{2}}\right)}.}$ (105)
Based on Eq. (105), we obtain
$\lim_{\gamma\to+\infty}\frac{Q^{-1}\left(\frac{\omega_{1}}{\sqrt{\Psi(\gamma)}}\right)}{\sqrt{\Psi(\gamma)}}$ $\leq\lim_{\gamma\to+\infty}\frac{\sqrt{W\left(\frac{\Psi(\gamma)}{2\pi\omega_{1}^{2}}\right)}}{\sqrt{\Psi(\gamma)}}$
$=\frac{1}{\sqrt{\log_{2}e}}\lim_{\gamma\to+\infty}\frac{\sqrt{\log_{2}\Psi(\gamma)-\log_{2}2\pi\omega_{1}^{2}}}{\sqrt{\Psi(\gamma)}}$ (106a)
$=\frac{1}{\sqrt{\log_{2}e}}\lim_{\gamma\to+\infty}\sqrt{\frac{\log_{2}\Psi(\gamma)}{\Psi(\gamma)}}$
$=0.$
Eq. (106a) holds because, according to [48], for $y>e$, the Lambert W function satisfies
${\color[rgb]{0,0,0}\ln y-\ln\ln y+\frac{\ln\ln y}{2\ln y}\leq W(y)\leq\ln y-\ln\ln y+\frac{e}{e-1}\frac{\ln\ln y}{\ln y}.}$ (107)
Thus, for $y>e$, we obtain
${\color[rgb]{0,0,0}\lim_{y\to+\infty}\frac{\sqrt{W(y)}}{\sqrt{\log_{2}y}}=\frac{1}{\sqrt{\log_{2}e}}.}$ (108)
Since $\frac{\Psi(\gamma)}{2\pi\omega_{1}^{2}}$ approaches +$\infty$+ ∞ with $\gamma\to+\infty$, we can apply Eq. (108) into Eq. (106a). The proof is completed.
Appendix D Proof of Theorem 5
Based on the definition of $s_{\infty}$ in Eq. (44) and $\tilde{R}^{\epsilon}_{N}[n]=R^{\epsilon}_{N}[n]-G(N,|h[n]|^{2}\gamma)$, we have
$s_{\infty}$ $=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\mathbb{E}_{|h|^{2}}\left\{e^{-T\theta R_{N}^{\epsilon}[n]}\right\}}{\log_{2}\gamma}$ (109)
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\mathbb{E}_{|h|^{2}}\left\{e^{-T\theta\left(R_{N}^{\epsilon}[n]-\tilde{R}_{N}^{\epsilon}[n]\right)}e^{-T\theta\tilde{R}_{N}^{\epsilon}[n]}\right\}}{\log_{2}\gamma}$
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\mathbb{E}_{|h|^{2}}\left\{e^{-T\theta\tilde{R}_{N}^{\epsilon}[n]}e^{-T\theta G(\Psi(\gamma),|h[n]|^{2}\gamma)}\right\}}{\log_{2}\gamma}.$
Since $N=\Psi(\gamma)$ and $N\theta=\varrho$, we have $\theta=\frac{\varrho}{\Psi(\gamma)}$. In Section II, we have proved that if $\Psi(\gamma)$ satisfies Theorem 2, Eq. (22) holds with a given $\epsilon$. Thus, we have
${\color[rgb]{0,0,0}\left|\theta G(N,|h[n]|^{2}\gamma)\right|=\left|\frac{\varrho G(N,|h[n]|^{2}\gamma)}{\Psi(\gamma)}\right|\leq\varrho\nu.}$ (110)
Thus, we have that $c_{5}\leq e^{-T\theta G(N,|h[n]|^{2}\gamma)}\leq c_{6}$, where $c_{i}$ is a positive constant, $i=5,6$. Thus, under the above conditions, we obtain
$s_{\infty}=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\mathbb{E}_{|h|^{2}}\left\{e^{-T\theta\tilde{R}_{N}^{\epsilon}[n]}\right\}}{\log_{2}\gamma}.$ (111)
According to Eq. (111), it is obvious that $s_{\infty}=1$ for the AWGN channel. For the Nakagami-$m$ fading channel, we have
$s_{\infty}$ $\!=\!-\frac{1}{\theta NT}\!\lim_{\gamma\to+\infty}\!\!\frac{\ln\!\int_{0}^{\infty}\!(1\!+\!\gamma x)^{-\theta NT\log_{2}e}e^{\theta T\sqrt{N\left(1-\frac{1}{(1+\gamma x)^{2}}\right)}Q^{-1}(\epsilon)\log_{2}e-\theta T\frac{\log_{2}N}{2}}\!\left(\frac{m}{\Omega}\right)^{m}\!\frac{x^{m-1}}{\Gamma(m)}e^{-\frac{m}{\Omega}x}{\rm d}x}{\log_{2}\gamma}$ (112)
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\int_{0}^{\infty}(1+\gamma x)^{-\theta NT\log_{2}e}e^{\theta T\sqrt{N\left(1-\frac{1}{(1+\gamma x)^{2}}\right)}Q^{-1}(\epsilon)\log_{2}e-\theta T\frac{\log_{2}N}{2}}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x}{\log_{2}\gamma}.$
We then divide the proof into four parts. First, we focus on the upper and lower bounds of $s_{\infty}$ for $m\geq 1$. Then we present the upper and lower bounds of $s_{\infty}$ for $m\in[0.5,1)$. For simplicity, we let $\beta=\sqrt{N}Q^{-1}(\epsilon)\log_{2}e$ in this paper.
D-A Upper bound of $s_{\infty}$ for $m\geq 1$
For $m\geq 1$, we first derive the upper bound of $s_{\infty}$. Since $R_{\epsilon}[n]\leq N\log_{2}\left(1+|h[n]|^{2}\gamma\right)$, we obtain
$s_{\infty}$ $\leq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[\int_{0}^{\infty}(1+\gamma x)^{-\theta NT\log_{2}e}\left(\frac{m}{\Omega}\right)^{m}e^{-\theta T\frac{\log_{2}N}{2}}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x\right]}{\log_{2}\gamma}$ (113)
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x\right]}{\log_{2}\gamma}.$
D-A1 $1<\frac{m}{\theta NT\log_{2}e}$
For $1<\frac{m}{\theta NT\log_{2}e}$, we find
$\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ $\leq\int_{0}^{\infty}x^{m-\theta NT\log_{2}e-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ (114)
$=\left(\frac{\Omega}{m}\right)^{m-\theta NT\log_{2}e}\int_{0}^{\infty}t^{m-\theta NT\log_{2}e-1}e^{-t}{\rm d}t.$
According to [14], $0<\int_{0}^{\infty}z^{a}e^{-x}{\rm d}x<\infty$ if $a>-1$. Since in this case $m>\theta NT\log_{2}e$, we obtain $m-\theta NT\log_{2}e-1>-1$. Therefore, $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ is bounded by a finite constant. Thus, under the condition $1<\frac{m}{\theta NT\log_{2}e}$, Eq. (113) can be further expressed as
$s_{\infty}\leq\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}=1.$ (115)
Eq. (115) holds because $\lim\limits_{\gamma\to+\infty}\frac{\log_{2}N}{N}=\lim\limits_{\gamma\to+\infty}\frac{\log_{2}\Psi(\gamma)}{2\Psi(\gamma)}=0$ holds under the condition that $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$.
D-A2 $1>\frac{m}{\theta NT\log_{2}e}$
For $1>\frac{m}{\theta NT\log_{2}e}$, we focus on the integral $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$. We find that
$\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$
$=$ $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e+m-1}\cdot\left(\frac{x}{\frac{1}{\gamma}+x}\right)^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$
$\geq$ $\int_{\frac{1}{\gamma}}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e+m-1}\left(\frac{1}{2}\right)^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ (116a)
$=$ $\left(\frac{1}{2}\right)^{m-1}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}e^{\frac{m}{\gamma\Omega}}\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right),$ (116b)
where Eq. (116a) holds because $\frac{x}{\frac{1}{\gamma}+x}$ an increasing function for $x\in\big{[}\frac{1}{\gamma},\infty\big{)}$.
By substituting Eq. (116b) into Eq. (113), we obtain
$s_{\infty}$ $\!\leq\!-\frac{1}{\theta NT}\!\!\lim_{\gamma\to+\infty}\!\!\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}\left(\frac{1}{2}\right)^{m-1}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}e^{\frac{m}{\gamma\Omega}}\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right)\right]}{\log_{2}\gamma}$
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-m}\left(\frac{1}{2}\right)^{\theta NT\log_{2}e-1}e^{\frac{m}{\gamma\Omega}}\frac{\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right)}{\left(\frac{2m}{\Omega\gamma}\right)^{m-\theta NT\log_{2}e}}\right]}{\log_{2}\gamma}$
$=\!-\frac{1}{\theta NT}\left(\!\!\lim_{\gamma\to+\infty}\!\!\frac{-m\ln\gamma}{\log_{2}\gamma}\!+\!\lim_{\gamma\to+\infty}\!\!\frac{\ln\left[\frac{\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right)}{\left(\frac{2m}{\Omega\gamma}\right)^{m-\theta NT\log_{2}e}}\right]}{\log_{2}\gamma}\!+\!\!\lim_{\gamma\to+\infty}\frac{m}{\Omega\gamma\log_{2}\gamma}\right)\!+\!\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}$
$=-\frac{1}{\theta NT}\left(\lim_{\gamma\to+\infty}\frac{-m\ln\gamma}{\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{m}{\Omega\gamma\log_{2}\gamma}\right)$ (117a)
$=\frac{m}{\theta NT\log_{2}e},$ (117b)
where Eq. (117a) holds because according to [49], for $s<0$,
$\lim_{x\to 0}\frac{\Gamma(s,x)}{x^{s}}=-\frac{1}{s}.$ (118)
D-A3 $1=\frac{m}{\theta NT\log_{2}e}$
For $1=\frac{m}{\theta NT\log_{2}e}$, $\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right)=E_{1}\left(\frac{2m}{\Omega\gamma}\right)$. By substituting $\Gamma\left(m-\theta NT\log_{2}e,\frac{2m}{\Omega\gamma}\right)=E_{1}\left(\frac{2m}{\Omega\gamma}\right)$ and Eq. (116b) into Eq. (113), we obtain
$s_{\infty}$ $\leq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}\left(\frac{1}{2}\right)^{m-1}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}e^{\frac{m}{\gamma\Omega}}E_{1}\left(\frac{2m}{\Omega\gamma}\right)\right]}{\log_{2}\gamma}$
$\leq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}\left(\frac{1}{2}\right)^{m}e^{-\frac{m}{\Omega\gamma}}\ln\left(1+\frac{\Omega\gamma}{m}\right)\right]}{\log_{2}\gamma}$ (119a)
$=\!-\!\frac{1}{\theta NT}\!\left(\!\lim_{\gamma\to+\infty}\!\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}\!-\!\!\lim_{\gamma\to+\infty}\!\frac{m}{\Omega\gamma\log_{2}\gamma}\!+\!\lim_{\gamma\to+\infty}\!\frac{\ln\ln\left(1+\frac{\Omega\gamma}{m}\right)}{\log_{2}\gamma}\!\right)\!+\!\lim_{\gamma\to+\infty}\!\frac{\log_{2}N}{2N}$
$=1.$
Eq. (119a) holds because according to [50]
$E_{1}(x)>\frac{1}{2}e^{-x}\ln\left(1+\frac{2}{x}\right),\quad x>0.\vspace{-2mm}$ (120)
Now we have completed the proof for the upper bound of $s_{\infty}$ for $m\geq 1$.
D-B Lower bound of $s_{\infty}$ for $m\geq 1$
Since $\sqrt{1-x}\leq 1$ with $x\in(0,1)$, the lower bound of $s_{\infty}$ in the Nakagami-$m$ fading channel is given by
$s_{\infty}$ $\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\int_{0}^{\infty}(1+\gamma x)^{-\theta NT\log_{2}e}e^{\theta T\beta-\theta T\frac{\log_{2}N}{2}}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x\right]}{\log_{2}\gamma}$
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x\right]}{\log_{2}\gamma}$ (121a)
$\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e+m-1}e^{-\frac{m}{\Omega}x}{\rm d}x\right]}{\log_{2}\gamma}$ (121b)
$=\!\!-\frac{1}{\theta NT}\!\!\lim_{\gamma\to+\infty}\!\!\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}e^{\frac{m}{\gamma\Omega}}\Gamma\left(m-\theta NT\log_{2}e,\frac{m}{\Omega\gamma}\right)\right]}{\log_{2}\gamma}.$
Eq. (121b) holds since $m\geq 1$ and $\frac{1}{\gamma}+x>x$ for $x>0$. Since $\frac{m}{\Omega}$ is a finite constant, $\lim\limits_{\gamma\to+\infty}\frac{m}{\Omega\gamma}=0$.
D-B1 $1<\frac{m}{\theta NT\log_{2}e}$
According to [51], for $s\in(0,1]\cup[2,+\infty)$, $\Gamma(s,x)$ satisfies
$\Gamma(s,x)<\frac{(x+v_{s})^{s}-x^{s}}{v_{s}s}e^{-x},$ (122)
where $v_{s}=\Gamma(s+1)^{\frac{1}{s-1}}>0$. For simplicity, we let $c_{7}=\Gamma(m+1-\theta NT\log_{2}e)^{\frac{1}{m-1-\theta NT\log_{2}e}}$. Based on Eq. (122), for $m-\theta NT\log_{2}e\in(0,1]\cup[2,\infty)$, we obtain
$\!s_{\infty}$ $\geq\!\!-\frac{1}{\theta NT}\!\!\lim_{\gamma\to+\infty}\!\!\!\frac{\ln\!\left[\!e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}\!\frac{\left(\frac{m}{\Omega\gamma}+c_{7}\right)^{m-\theta NT\log_{2}e}-\left(\frac{m}{\Omega\gamma}\right)^{m-\theta NT\log_{2}e}}{c_{7}(m-\theta NT\log_{2}e)}\right]}{\log_{2}\gamma}$
$=\!\!-\frac{1}{\theta NT}\!\left(\!\!\lim_{\gamma\to+\infty}\!\!\!\frac{\ln\left[\gamma^{-m}e^{\theta T\beta}\left((1+\frac{c_{7}\Omega}{m}\gamma)^{m-\theta NT\log_{2}e}-1\right)\right]}{\log_{2}\gamma}\!-\!\!\lim_{\gamma\to+\infty}\!\!\frac{\ln\left[c_{7}\left(m-\theta NT\log_{2}e\right)\right]}{\log_{2}\gamma}\right)$
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-m\ln\gamma}{\log_{2}\gamma}-\lim_{\gamma\to+\infty}\frac{\beta}{N\log_{2}\gamma}-\frac{m-\theta NT\log_{2}e}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left(1+\frac{c_{7}\Omega}{m}\gamma\right)}{\log_{2}\gamma}$ (123a)
$=1.$ (123b)
Eq. (123a) holds because $\lim\limits_{\gamma\to+\infty}\frac{\ln\left[\left(1+\frac{c_{7}\Omega}{m}\gamma\right)^{m-\theta NT\log_{2}e}-1\right]}{(m-\theta NT\log_{2}e)\ln\left(1+\frac{c_{7}\Omega}{m}\gamma\right)}=1$ with $c_{7}>0$ and $m-\theta NT\log_{2}e>0$. Eq. (123b) holds because $\lim\limits_{\gamma\to+\infty}\frac{\ln\left(1+\frac{c_{7}\Omega}{m}\gamma\right)}{\ln\gamma}=1$.
For $m-\theta NT\log_{2}e\in(1,2)$, we focus on the integral in Eq. (121a). When $m-\theta NT\log_{2}e\in(1,2)$, $\theta NT\log_{2}e<m$ holds. Thus, we find that $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ is upper bounded by a finite constant according to Eq. (114). Thus, we obtain that
$s_{\infty}\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}-\lim_{\gamma\to+\infty}\frac{\beta}{N\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}=1.$ (124)
D-B2 $1\geq\frac{m}{\theta NT\log_{2}e}$
For $m-\theta NT\log_{2}e<0$, we have
$s_{\infty}$ $\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-m}e^{\theta T\beta}e^{\frac{m}{\Omega\gamma}}\frac{\Gamma\left(m-\theta NT\log_{2}e,\frac{m}{\Omega\gamma}\right)}{\left(\frac{m}{\Omega\gamma}\right)^{m-\theta NT\log_{2}e}}\right]}{\log_{2}\gamma}$
$=-\frac{1}{\theta NT}\left(-\lim_{\gamma\to+\infty}\frac{m\ln\gamma}{\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{\theta T\beta}{\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{m}{\Omega\gamma\log_{2}\gamma}\right)+\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}$ (125a)
$=\frac{m}{\theta NT\log_{2}e},$
where Eq. (125a) holds according to Eq. (118).
D-B3 $1=\frac{m}{\theta NT\log_{2}e}$
For $m-\theta NT\log_{2}e=0$, $\Gamma\left(m-\theta NT\log_{2}e,\frac{m}{\Omega\gamma}\right)=E_{1}\left(\frac{m}{\Omega\gamma}\right)$. According to Eq. (121), we have
$s_{\infty}$ $\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}e^{\frac{m}{\gamma\Omega}}E_{1}\left(\frac{m}{\Omega\gamma}\right)\right]}{\log_{2}\gamma}$
$\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}\ln\left(1+\frac{\Omega\gamma}{m}\right)\right]}{\log_{2}\gamma}$ (126a)
$=-\frac{1}{\theta NT}\left(\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{\theta T\beta}{\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{\ln\ln\left(1+\frac{\Omega\gamma}{m}\right)}{\log_{2}\gamma}\right)\!+\!\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}$ (126b)
$=1.$
Eq. (126a) holds because according to [50]
$E_{1}(x)<e^{-x}\ln\left(1+\frac{1}{x}\right),\quad x>0.$ (127)
Therefore, according to the squeeze theorem, we have completed the proof for $m\geq 1$ based on the results in Eqs. (115), (117), (119), (123), (125), and (126).
D-C Upper bound of $s_{\infty}$ for $m\in[0.5,1)$
For $10.5\leq m<$ 1, there are some differences in the proof, which mainly lie in the proof of the upper bound with $1>\frac{m}{\theta NT\log_{2}e}$ and the lower bound.
For the upper bound with $1\leq\frac{m}{\theta NT\log_{2}e}$, the proof is same with Eqs. (115) and (119). For the upper bound with $1>\frac{m}{\theta NT\log_{2}e}$, Eq. (116) is changed into
$\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$
$=$ $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e+m-1}\left(\frac{1}{\gamma x}+1\right)^{1-m}e^{-\frac{m}{\Omega}x}{\rm d}x$
$\geq$ $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e+m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ (128a)
$=$ $\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}e^{\frac{m}{\Omega\gamma}}\Gamma\left(m-\theta NT\log_{2}e,\frac{m}{\Omega\gamma}\right).$
Eq. (128a) holds because $01-m>$ 0 and $11+\frac{1}{\gamma x}>$ 1 for $x>0$.
Then, by substituting Eq. (128) into Eq. (113), we obtain
$s_{\infty}$ $\leq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}\left(\frac{m}{\Omega}\right)^{\theta NT\log_{2}e-m}e^{\frac{m}{\Omega\gamma}}\Gamma\left(m-\theta NT\log_{2}e,\frac{m}{\Omega\gamma}\right)\right]}{\log_{2}\gamma}$
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-m}e^{\frac{m}{\Omega\gamma}}\frac{\Gamma\left(m-\theta NT\log_{2}e,\frac{m}{\Omega\gamma}\right)}{\left(\frac{m}{\Omega\gamma}\right)^{m-\theta NT\log_{2}e}}\right]}{\log_{2}\gamma}$
$=-\frac{1}{\theta NT}\left(\lim_{\gamma\to+\infty}\frac{-m\ln\gamma}{\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{m}{\Omega\gamma\log_{2}\gamma}\right)+\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}$ (129a)
$=\frac{m}{\theta NT\log_{2}e}.$
Eq. (129a) holds because $m-\theta NT\log_{2}e<0$ so that Eq. (118) holds.
D-D Lower Bound of $s_{\infty}$ for $m\in[0.5,1)$
Then, we focus on the lower bound of $s_{\infty}$ with $10.5\leq m<$ 1. The essence of this part of the proof is still the scaling of $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$. According to Eq. (121), the lower bound of $s_{\infty}$ can be expressed as
$s_{\infty}\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x\right]}{\log_{2}\gamma}.$ (130)
D-D1 $1<\frac{m}{\theta NT\log_{2}e}$
For $1<\frac{m}{\theta NT\log_{2}e}$, we transform Eq. (130) into
$s_{\infty}\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\int_{0}^{\infty}x^{-\theta NT\log_{2}e+m-1}e^{-\frac{m}{\Omega}x}{\rm d}x\right]}{\log_{2}\gamma}.$ (131)
Eq. (131) holds because $-\theta NT\log_{2}e<0$ and $\frac{1}{\gamma}+x>x$. As we mentioned before, $\int_{0}^{\infty}z^{a}e^{-x}{\rm d}x<\infty$ holds when $a>-1$. Since in this case $1<\frac{m}{\theta NT\log_{2}e}$, $-\theta NT\log_{2}e+m-1>-1$. Thus, in this case $\int_{0}^{\infty}x^{-\theta NT\log_{2}e+m-1}e^{-\frac{m}{\Omega}x}{\rm d}x=\left(\frac{\Omega}{m}\right)^{m-\theta NT\log_{2}e}\int_{0}^{\infty}t^{m-\theta NT\log_{2}e-1}e^{-t}{\rm d}t$ is a finite constant. Then, we obtain
$s_{\infty}\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}-\frac{1}{N}\lim_{\gamma\to+\infty}\frac{\beta}{\log_{2}\gamma}+\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}=1.$ (132)
D-D2 $1\geq\frac{m}{\theta NT\log_{2}e}$
For $1\geq\frac{m}{\theta NT\log_{2}e}$, we find
$\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$
$=\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e+m}\left(\frac{x}{\frac{1}{\gamma}+x}\right)^{m-\frac{1}{2}}\frac{x^{-\frac{1}{2}}}{\left(\frac{1}{\gamma}+x\right)^{\frac{1}{2}}}e^{-\frac{m}{\Omega}x}{\rm d}x$
$\leq\left(\frac{1}{\gamma}\right)^{-\theta NT\log_{2}e+m}\int_{0}^{\infty}\frac{e^{-\frac{m}{\Omega}x}}{\sqrt{x\left(\frac{1}{\gamma}+x\right)}}{\rm d}x$ (133a)
$=\gamma^{\theta NT\log_{2}e-m}e^{\frac{m}{2\Omega\gamma}}K_{0}\left(\frac{m}{2\Omega\gamma}\right),$
where $K_{v}(z)$ is the modified Bessel function of the second kind [50]. Eq. (133a) holds because in this case $m-\frac{1}{2}\geq 0$ and $\frac{x}{\frac{1}{\gamma}+x}\leq 1$. According to [50], for the modified Bessel function of the second kind $K_{v}(z)$ with $v=0$, we have
$\lim_{z\to 0}\frac{K_{0}(z)}{-\ln(z)}=c_{8},$ (134)
where $c_{8}>0$ is a constant. Based on this conclusion, we transformed Eq. (133) into
$\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x\leq\gamma^{\theta NT\log_{2}e-m}e^{\frac{m}{2\Omega\gamma}}\ln\left(\frac{2\Omega\gamma}{m}\right)\frac{K_{0}\left(\frac{m}{2\Omega\gamma}\right)}{-\ln\left(\frac{m}{2\Omega\gamma}\right)}.$ (135)
By substituting Eq. (135) into Eq. (130), we obtain
$s_{\infty}$ $\geq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[e^{-\theta T\frac{\log_{2}N}{2}}\gamma^{-m}e^{\theta T\beta}e^{\frac{m}{2\Omega\gamma}}\ln\left(\frac{2\Omega\gamma}{m}\right)\frac{K_{0}\left(\frac{m}{2\Omega\gamma}\right)}{-\ln\left(\frac{m}{2\Omega\gamma}\right)}\right]}{\log_{2}\gamma}$
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\left(\frac{-m\ln\gamma}{\log_{2}\gamma}+\frac{\theta T\beta}{\log_{2}\gamma}+\frac{\ln\ln\gamma}{\log_{2}\gamma}+\frac{\ln\frac{K_{0}\left(\frac{m}{2\Omega\gamma}\right)}{-\ln\left(\frac{m}{2\Omega\gamma}\right)}}{\log_{2}\gamma}\right)+\lim_{\gamma\to+\infty}\frac{\log_{2}N}{2N}$ (136a)
$=\frac{m}{\theta NT\log_{2}e},$ (136b)
where Eq. (136b) holds according to Eq. (134). According to the squeeze theorem, we have completed the proof for $m\in[0.5,1)$. Now we have completed the proof.
Appendix E Proof of Theorem 6
With the $\Psi(\gamma)$ satisfies Theorem 3, Eq. (111) still holds. For the AWGN channel, we can easily obtain
$s_{\infty}$ $=\lim_{\gamma\to+\infty}\frac{\tilde{R}_{N}^{\epsilon}}{N\log_{2}\gamma}$ (137a)
$=1-\log_{2}e\lim_{\gamma\to+\infty}\frac{Q^{-1}\left(\epsilon(\gamma)\right)}{\sqrt{N}\log_{2}\gamma}.$
For an FBL-SISO system in a Nakagami-$m$ fading channel, we will prove that the following two inequalities hold.
$\lim_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}+\log_{2}e\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}\geq\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\}.$ (138)
$\lim_{\gamma\to+\infty}\frac{\Lambda(\gamma)}{\log_{2}\gamma}+(1-c_{3})\log_{2}e\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}\leq\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\}.$ (139)
In Eq. (139), $c_{3}>0$ is an arbitrary constant. Since $c_{3}>0$ can be an arbitrarily small constant, we obtain Eq. (49) based on Eqs. (138) and (139).
The derivation of Eq. (138) is based on the derivation of the lower bounds of $s_{\infty}$ in the proof of Theorem 5. For $m\geq 1$, we can directly obtain Eq. (138) according to Eqs. (123a), (125a), and (126b). For $m\in[0.5,1)$, we can obtain Eq. (138) according to Eqs. (132) and (136a). Thus, we omit the details here.
The derivation of Eq. (139) is based on the derivation of the upper bounds of $s_{\infty}$. Different from the proof of Theorem 5, we start the proof with the help of $\sqrt{1-x}\geq 1-\sqrt{x}$ with $x\in(0,1)$. We have
$s_{\infty}$ $\leq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{e^{-\theta T\frac{\log_{2}N}{2}}\ln\left[\int_{0}^{\infty}e^{-\theta T\left(N\log_{2}(1+\gamma x)-\beta\left(1-\frac{1}{1+\gamma x}\right)\right)}\left(\frac{m}{\Omega}\right)^{m}\frac{x^{m-1}}{\Gamma(m)}e^{-\frac{m}{\Omega}x}{\rm d}x\right]}{\log_{2}\gamma}$ (140)
$=-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}x^{m-1}e^{-\left(\frac{m}{\Omega}x+\frac{\theta T\beta}{1+\gamma x}\right)}{\rm d}x\right]}{\log_{2}\gamma}.$
Based on $\frac{x}{x+1}\leq\ln(1+x)$ for $x>0$, we have $\frac{1}{1+\gamma x}\leq\ln\left(1+\frac{1}{\gamma x}\right)$. Thus, we transform Eq. (140) into
$s_{\infty}\leq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}\left(1+\frac{1}{\gamma x}\right)^{-\theta T\beta}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x\right]}{\log_{2}\gamma}.$ (141)
The integral in the upper bound of $s_{\infty}$ is given by
$\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e}\left(1+\frac{1}{\gamma x}\right)^{-\theta T\beta}x^{m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$ (142)
$=$ $\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e+m-1}\left(\frac{1}{\gamma x}+1\right)^{1-m-\theta T\beta}e^{-\frac{m}{\Omega}x}{\rm d}x.$
Note that, with $\gamma\to+\infty$, $\beta\to+\infty$. Thus, $01-m-\theta T\beta<$ 0. We obtain
$\int_{0}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e+m-1}\left(\frac{1}{\gamma x}+1\right)^{1-m-\theta T\beta}e^{-\frac{m}{\Omega}x}{\rm d}x$ (143)
$\geq$ $\left(\frac{c_{9}}{c_{9}+1}\right)^{\theta T\beta+m-1}\int_{\frac{c_{9}}{\gamma}}^{\infty}\left(\frac{1}{\gamma}+x\right)^{-\theta NT\log_{2}e+m-1}e^{-\frac{m}{\Omega}x}{\rm d}x$
$=$ $\left(\frac{c_{9}}{c_{9}+1}\right)^{\theta T\beta+m-1}\left(\frac{m}{\Omega}\right)^{\theta N\log_{2}e-m}e^{\frac{m}{\Omega\gamma}}\Gamma\left(m-\theta NT\log_{2}e,\frac{m(1+c_{9})}{\Omega\gamma}\right),$
where $c_{9}$ is an arbitrarily positive constant.
Based on this result, we obtain
$s_{\infty}\!\leq\!\!-\frac{1}{\theta NT}\!\!\lim_{\gamma\to+\infty}\!\!\!\frac{\ln\!\left[\!\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\left(\!\frac{c_{9}}{c_{9}+1}\!\right)^{\theta T\beta+m-1}\!\!\left(\frac{m}{\Omega}\right)^{\theta N\log_{2}e-m}\!e^{\frac{m}{\Omega\gamma}}\Gamma\left(\!m\!-\!\theta NT\log_{2}e,\frac{m(1+c_{9})}{\Omega\gamma}\!\right)\!\right]}{\log_{2}\gamma}.$ (144)
For $1>\frac{m}{\theta NT\log_{2}e}$, by combining Eq. (118) with Eq. (144), we have
$s_{\infty}$ $\leq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{-m\ln\gamma}{\log_{2}\gamma}-\left(1+\ln\frac{c_{9}}{c_{9}+1}\right)\lim_{\gamma\to+\infty}\frac{\beta}{N\log_{2}\gamma}$ (145)
$=\frac{m}{\theta NT\log_{2}e}-\log_{2}e\left(1-\ln\left(1+\frac{1}{c_{9}}\right)\right)\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}.$
Note that, since $c_{9}$ is an arbitrarily positive constant, $\ln\left(1+\frac{1}{c_{9}}\right)$ is an arbitrarily positive constant.
For $1\leq\frac{m}{\theta NT\log_{2}e}$, we will divide the proof into two parts.
The first part is $1\leq\frac{m}{\theta NT\log_{2}e}\leq 1+\frac{1}{\theta NT\log_{2}e}$ and $\frac{m}{\theta NT\log_{2}e}\geq 1+\frac{2}{\theta NT\log_{2}e}$, which means $0\leq m-\theta NT\log_{2}e\leq 1$ and $m-\theta NT\log_{2}e\geq 2$. We will utilize the conclusion that $\Gamma(s,x)\geq e^{-x}(1+x)^{s-1}$ for $x>0$ with $s\leq 1$ and $s\geq 2$. This can be easily proved by Jensen’s inequality. $(t+x)^{s-1}$ is a convex function of $t$ with a $s\leq 1$ and $s\geq 2$. Thus, we have
$\int_{x}^{\infty}y^{s-1}e^{-y}{\rm d}y=e^{-x}\int_{0}^{\infty}(t+x)^{s-1}e^{-t}{\rm d}t\geq e^{-x}\left(\int_{0}^{\infty}(t+x)e^{-t}{\rm d}t\right)^{s-1}=e^{-x}(1+x)^{s-1}.$ (146)
Now we have proved that $\Gamma(s,x)\geq e^{-x}(1+x)^{s-1}$ holds for $s\leq 1$ and $s\geq 2$.
Thus, by substituting Eq. (146) into Eq. (144), we have
$s_{\infty}$ $\leq-\frac{1}{\theta NT}\lim_{\gamma\to+\infty}\frac{\ln\left[\gamma^{-\theta NT\log_{2}e}e^{\theta T\beta}\left(\frac{c_{9}}{c_{9}+1}\right)^{\theta T\beta}e^{-\frac{c_{9}m}{\Omega\gamma}}\left(1+\frac{m(1+c_{9})}{\Omega\gamma}\right)^{m-\theta NT\log_{2}e-1}\right]}{\log_{2}\gamma}$ (147)
$=\!-\frac{1}{\theta NT}\!\!\lim_{\gamma\to+\infty}\!\!\left(\!\frac{-\theta NT\log_{2}e\ln\gamma}{\log_{2}\gamma}\!+\!\left(\!1+\ln\frac{c_{9}}{c_{9}+1}\!\right)\frac{\theta T\beta}{\log_{2}\gamma}\!+\!(m\!-\!\theta NT\!-\!1)\log_{2}e\frac{m(1+c_{9})}{\gamma\log_{2}\gamma}\!\right)$
$=1-\log_{2}e\left(1-\ln\left(1+\frac{1}{c_{9}}\right)\right)\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}.$
The first part is $1+\frac{1}{\theta NT\log_{2}e}<\frac{m}{\theta NT\log_{2}e}<1+\frac{2}{\theta NT\log_{2}e}$, which means $21<m-\theta NT\log_{2}e<$ 2. For $\Gamma(s,x)$ with $21<s<$ 2, according to [51], we have
$\Gamma(s,x)>\frac{(x+v_{s})^{s}-x^{s}}{v_{s}s}e^{-x},$ (148)
where $v_{s}$ is defined after Eq. (122). By substituting Eq. (148) into Eq. (144), we can obtain Eq. (149) with the similar procedures as shown in Eq. (123).
$s_{\infty}$ $\!\!\leq\!-\frac{1}{\theta NT}\!\lim_{\gamma\to+\infty}\!\!\frac{-m\ln\gamma}{\log_{2}\gamma}\!-\!\left(\!1\!+\!\ln\frac{c_{2}}{c_{2}+1}\right)\!\!\lim_{\gamma\to+\infty}\!\frac{\beta}{N\log_{2}\gamma}\!-\!\left(\frac{m}{\theta NT}\!-\!1\right)\!\lim_{\gamma\to+\infty}\!\!\frac{\ln\left(1+\frac{c_{1}\Omega(1+c_{2})}{m}\gamma\right)}{\log_{2}\gamma}$ (149)
$=1-\log_{2}e\left(1-\ln\left(1+\frac{1}{c_{2}}\right)\right)\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}.$
Now we have completed the proof of Eq. (139).
Appendix F Proof of Lemma 1
Based on the definition of $\varpi$, we obtain
$\lim_{\gamma\to+\infty}\frac{1}{\sqrt{N}}\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\partial\log_{2}\gamma}=\varpi\sqrt{\frac{1}{2\log_{2}e}}.$ (150)
Under the assumption that $\varpi$ is finite, we have
$\lim_{\gamma\to+\infty}\frac{1}{\sqrt{N}}\sqrt{\frac{-\log_{2}\epsilon(\gamma)}{\log_{2}^{2}\gamma}}$ $=\lim_{\gamma\to+\infty}\frac{\sqrt{-\log_{2}\epsilon(\gamma)}}{\sqrt{N}\log_{2}\gamma}$ (151a)
$=\lim_{\gamma\to+\infty}\frac{\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\partial\log_{2}\gamma}}{\sqrt{N}+\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\log_{2}\gamma}$ (151b)
$=\varpi\sqrt{\frac{1}{2\log_{2}e}}\lim_{\gamma\to+\infty}\frac{1}{1+\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{N}}}.$
Eq. (151b) holds according to L’Hospital’s rule. As defined in Lemma 1, we let $c_{4}$ denote $\lim_{\gamma\to+\infty}\frac{1}{1+\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{N}}}$. Since $\lim\limits_{\gamma\to+\infty}\Psi(\gamma)=+\infty$, $\frac{\partial\sqrt{N}}{\partial\log_{2}\gamma}\frac{\log_{2}\gamma}{\sqrt{N}}\geq 0$. Thus, $c_{3}\in[0,1]$.
Then, we let $Q^{-1}(\epsilon(\gamma))=x$, which can also be written as $\epsilon(\gamma)=Q(x)$. According to Eq. (102), we have
$x^{2}+\ln x^{2}\leq-2\ln\left(\sqrt{2\pi}\epsilon(\gamma)\right).$ (152)
By replacing $v$ by $\epsilon(\gamma)$ in Eq. (104), we have
${\color[rgb]{0,0,0}-\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}\leq x\leq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}.}$ (153)
Since $x>0$, we have $0<x\leq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}$. Similarly, according to Eq. (102), we obtain
$\left(x+\frac{1}{x}\right)^{2}+\ln\left(x+\frac{1}{x}\right)^{2}>x^{2}+\ln\left(x+\frac{1}{x}\right)^{2}\geq-2\ln\left(\sqrt{2\pi}\epsilon(\gamma)\right).$ (154)
Based on $W(\cdot)$, we obtain
${\color[rgb]{0,0,0}x+\frac{1}{x}\geq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}\text{ or }x+\frac{1}{x}\leq-\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}}$ (155)
Since $x>0$, we will utilize $x+\frac{1}{x}\geq\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}$ in the following.
Based on Eq. (153), we obtain
$\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}$ $\leq\lim_{\gamma\to+\infty}\frac{\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}}{\sqrt{N}\log_{2}\gamma}$
$=\lim_{\gamma\to+\infty}\frac{\sqrt{-2\log_{2}\epsilon(\gamma)}}{\sqrt{N}\log_{2}\gamma}\cdot\frac{\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}}{\sqrt{-2\log_{2}\epsilon(\gamma)}}$
$=\frac{1}{\sqrt{\log_{2}e}}\lim_{\gamma\to+\infty}\sqrt{\frac{-2\log_{2}\epsilon(\gamma)}{\log_{2}^{2}\gamma}}\cdot\frac{1}{\sqrt{N}}$ (156a)
$=\frac{c_{4}}{\log_{2}e}\varpi.$ (156b)
Eq. (156a) holds according to Eq. (108). Besides, Eq. (156b) holds according to Eq. (151).
Similarly, based on Eq. (155), we obtain
$\lim_{\gamma\to+\infty}\frac{Q^{-1}(\epsilon(\gamma))}{\sqrt{N}\log_{2}\gamma}\geq\lim_{\gamma\to+\infty}\frac{\sqrt{W\left(\frac{\epsilon^{-2}(\gamma)}{2\pi}\right)}}{\sqrt{N}\log_{2}\gamma}-\lim_{\gamma\to+\infty}\frac{1}{Q^{-1}(\epsilon(\gamma))\cdot\sqrt{N}\log_{2}\gamma}=\frac{c_{4}}{\log_{2}e}\varpi.$ (157)
Eq. (157) holds because with $\gamma\to+\infty$, $Q^{-1}(\epsilon(\gamma))\to+\infty$. Thus, according to the squeeze theorem, we have completed the proof.
Appendix G Proof of Corollary 3
Based on Lemma 1 in [52], we find that $\alpha_{A}(\theta)$ is an increasing function of $\theta$. Thus, the inverse function of $\alpha_{A}(\cdot)$ exists, which we denoted by $\alpha^{-1}_{A}(\cdot)$. According to Eq. (20), the $\theta$ for the AWGN channel can be denoted by
$\theta=\alpha_{A}^{-1}(R_{N}^{\epsilon}).$ (158)
Based on Eq. (19), we have $\chi\leq e^{-\theta L}$. By substituting Eq. (158) into $\chi\leq e^{-\theta L}\leq\chi_{\rm th}$, we have
$\alpha_{A}^{-1}(R_{N}^{\epsilon})\geq-\frac{1}{L}\ln\chi_{\rm th}.$ (159)
Since $\alpha_{A}(\theta)$ is an increasing function of $\theta$, we have $R_{N}^{\epsilon}\geq\alpha_{A}\left(-\frac{1}{L}\ln\chi\right)$. By substituting Eq. (7) into this inequality, we have
${\color[rgb]{0,0,0}NC-\sqrt{NV}Q^{-1}(\epsilon)+\frac{\log_{2}N}{2}+G(N,\gamma,\epsilon)\geq\alpha_{A}\left(-\frac{1}{L}\ln\chi\right).}$ (160)
Based on Eq. (160), we obtain Eq. (70).
Appendix H Proof of Lemma 2
We put our focus on the expectation term in the definition of EC. We define $g(x)$ as $-r(x)$. The derivative of $g(x)$ is given by
$g^{\prime}(x)=\frac{-(\ln 2)^{-1}\left(\gamma(1+\gamma x)\sqrt{(1+\gamma x)^{2}-1}-\sqrt{\frac{1}{N}}Q^{-1}(\epsilon)\gamma\right)}{(1+\gamma x)^{3}\sqrt{\big{(}1-(1+\gamma x)^{-2}\big{)}}}.$ (161)
$(1+\gamma x)\sqrt{(1+\gamma x)^{2}-1}$ is an increasing function of $x$, which is equals to 0 with $x=0$. Thus, based on Eq. (161), we find that there exists a $x^{*}$, when $x=x^{*}$, $g^{\prime}(x)=0$, $x<x^{*}$, $g^{\prime}(x)>0$, and $x>x^{*}$, $g^{\prime}(x)<0$. $x^{*}$ satisfies
$(1+\gamma x^{*})\sqrt{(1+\gamma x^{*})^{2}-1}=\frac{Q^{-1}(\epsilon)}{\sqrt{N}}.$ (162)
$x^{*}$ can be solved from Eq. (162) as
$x^{*}=\frac{1}{\gamma}\left(\sqrt{\frac{1+\sqrt{1+4\ell}}{2}}-1\right),$ (163)
which is an interior point of $(0,\infty)$. Then, according to Laplace’s method [53], the integral in the definition of EC can be expressed as
[…]
By substituting Eq. (164) into Eq. (16), we obtain Eq. (71).
Appendix I Proof of Theorem 8
Similar to the proof of Lemma 2, the derivative of $g(x)$ is given by
$g^{\prime}(x)=\frac{-(\ln 2)^{-1}\left(\Xi+\Xi^{\prime}x\right)\left((1+\Xi x)\sqrt{(1+\Xi x)^{2}-1}-\sqrt{\frac{1}{N}}Q^{-1}(\epsilon)\right)}{(1+\Xi x)^{3}\sqrt{\big{(}1-(1+\Xi x)^{-2}\big{)}}},$ (165)
where for simplicity we omit the argument of $\Xi$. By comparing Eqs. (165) and (161), we find that if $\Xi+\Xi^{\prime}x\geq 0$, the remaining analysis is identical with the proof of Lemma 2.
Appendix J Proof of Corollary 4
According to [28], the EC with the simple ARQ mechanism that has an unlimited number of retransmissions is given by
$\alpha_{S}(\theta)$ $=-\frac{1}{\theta T}\ln\mathbb{E}_{x}\left\{\epsilon+(1-\epsilon)e^{-\theta NTr(x)}\right\}$ (166)
$=-\frac{1}{\theta T}\ln\left(\epsilon+(1-\epsilon)\int_{0}^{\infty}e^{-\theta NTr(x)}f(x){\rm d}x\right).$
Similar to the proofs of Lemma 2 and Theorem 8, if the power allocation scheme satisfies $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$, the integral $\int_{0}^{\infty}e^{-\theta NTr(x)}f(x){\rm d}x$ can be approximated as shown in Eq. (164) with the help of Laplace’s method. By substituting Eq. (164) into Eq. (166), we obtain Eq. (78).